### v1

# A Proxy-Free Strategy for Practically Improving the Poisoning Efficiency in Backdoor Attacks

Abstract
Poisoning efficiency is a crucial factor in poisoning-based backdoor attacks. Attackers prefer to use as few poisoned samples as possible to achieve the same level of attack strength, in order to remain undetected. Efficient triggers have significantly improved poisoning efficiency, but there is still room for improvement. Recently, selecting efficient samples has shown promise, but it requires a proxy backdoor injection task to find an efficient poisoned sample set, which can lead to performance degradation if the proxy attack settings are different from the actual settings used by the victims. In this paper, we propose a novel Proxy-Free Strategy (PFS) that selects efficient poisoned samples based on individual similarity and set diversity, effectively addressing this issue. We evaluate the proposed strategy on several datasets, triggers, poisoning ratios, architectures, and training hyperparameters. Our experimental results demonstrate that PFS achieves higher backdoor attack strength while 500 faster than previous proxy-based selection approaches.
1 Introduction
The abundance of training data plays a vital role in the success of Deep Neural Networks (DNNs). For instance, GPT-3 [2], a deep learning model with 175 billion parameters, owes its effectiveness in various language processing tasks to its pretraining on a massive corpus of 45 TB of text data. With the increasing demand for data, users and businesses are turning to third-party sources or online collection as easier means of obtaining it. However, recent studies [4, 12, 3, 11] have demonstrated that such practices can be maliciously exploited by attackers to contaminate training data, thereby negatively impacting the functionality and reliability of trained models.
During the training phase, one of the significant threats is called backdoor attacks [4, 12, 27]. This attack involves injecting a covert backdoor into the DNN by introducing a small number of poisoned samples into the benign training set. Although the model appears normal on benign inputs, a predefined trigger can activate the infected model, forcing its predictions to align with the attacker’s objective. As research continues to reveal this threat in various tasks such as speaker verification [46] and malware detection [20], attention to backdoor attacks is increasing.
Poisoning efficiency is a crucial indicator for attackers attempting to launch data poisoning attacks on DNNs [47, 22, 39, 40]. The aim is to contaminate the training dataset with as few poisoned samples as possible while achieving the desired outcome. A perfect scenario is if a backdoor can be successfully implanted by poisoning only one sample in the dataset, significantly enhancing the attack’s stealthiness. Currently, methods to improve poisoning efficiency fall into two categories: designing efficient triggers [47, 48, 44] and selecting efficient samples [39]. The former aims to find trigger patterns that are easier for models to learn, while the latter focuses on selecting and representative poisoned samples. This paper is concerned with the latter approach.
A recent study [39] on sample selection has yielded significant results, demonstrating that the same attack success rate can be achieved with only to of the poisoned samples compared to random selection strategies. However, we have found that this method requires a proxy backdoor injection task to evaluate the contribution of each poisoned sample, resulting in a constructed poisoned set closely related to the attacker’s specific proxy attack. This presents a challenge, as the effectiveness of the selected set will be damaged when the attacker’s proxy attack settings differ from the actual settings of the victims, which is a common scenario in practice. Moreover, conducting the proxy task significantly increases the execution time of the algorithm, limiting its scalability on large datasets.
In this paper, we address the above issue and present a Proxy-Free Strategy (PFS) that acquires higher backdoor attack strength while 500 faster than previous selection method. Our main contributions are as follows:
We conduct an analysis of the forensic features of efficient poisoned samples in backdoor attacks, with a specific focus on two key factors: the similarity between clean and corresponding poisoned samples, and the diversity of the poisoned sample set.
Based on this analysis, we propose a novel selection method called the Proxy-Free Strategy (PFS), which does not require a proxy attack task, thereby addressing the issue of inconsistent settings and significantly reducing the time consumption.
we evaluate PFS on three datasets: CIFAR-10, CIFAR-100, and Tiny-ImageNet, and our results consistently demonstrate its effectiveness. Notably, PFS achieves comparable or even higher attack success rates while 500 faster than previous proxy-based selection approaches, making it more scalable for larger datasets or complex models.
2 Related Works
2.1 Backdoor Attacks
Backdoor attacks aim to embed hidden Trojans into DNNs, enabling the attacked models to perform well on clean samples while misbehaving on samples with specifically designed triggers. These attacks can occur at any stage of DNN development, including code [1, 33], outsourcing, pretrained [36, 10, 19], data collection [12, 4, 35, 26], collaborative learning [42, 31], and post-deployment [32, 17]. Among these, poisoning-based backdoor attacks, which involve injecting backdoor triggers by modifying the training data, are the most straightforward and widely adopted. Recent studies have focused on improving poisoning efficiency, which can be divided into two categories.
Designing Efficient Trigger. Trigger design is a crucial aspect of backdoor attacks. Chen et al. [4] proposed a blended strategy that blends clean samples with the trigger to evade human detection. Later, studies aimed to create more efficient and invisible triggers by adopting natural patterns such as warping [30], rotation [38], style transfer [5], frequency [8, 45], and reflection [28]. In some studies, researchers have taken inspiration from Universal Adversarial Perturbations (UAPs) [29] and optimized an UAP on a pre-trained clean model to create a trigger, which is an effective and widely used approach [48, 21, 6]. Li et al. [23] proposed a different approach where they use GANs models to acquire a sample-specific trigger, which is a departure from previous works that adopt universal triggers.
Selecting Efficient Poisoning Samples. Efficient sample selection for poisoning attacks is a crucial but under-explored area, and is separate from the trigger design. Xia et al. [39] were among the first to explore the contribution of different data to backdoor injection, discovering that not all poisoned samples contribute equally and that appropriate sample selection can significantly improve the efficiency of data in backdoor attacks. However, their study [39] relies on proxy attacks to identify the efficient subset, which suffers from performance degradation when gaps exist (e.g., in Transform (F) and Hyperparameters (H) as shown in Figure 1 between the proxy poisoning attack and the actual poisoning process used by victims. Therefore, a sample selection strategy that is not based on proxy attacks is urgently needed, and is the main contribution of our paper.
2.2 Backdoor Defenses
Backdoor attacks pose a significant threat to the security of DNN models, and several defense methods have been proposed to mitigate these threats. Existing backdoor defenses can be broadly classified into two categories: backdoor detection and backdoor erasing. Backdoor detection aims to detect poisoning in input data or DNN models used [9, 14, 7, 41]. While these methods have demonstrated promising detection precision, they cannot completely mitigate the potential threats of backdoor models. Backdoor erasing aims to purify the threatened models by erasing the backdoor during or post training. During training, clean models can be trained on poisoned data, as proposed by Li et al. [24]. Post-training, models can be fine-tuned, distilled, or pruned to erase the backdoor from the attacked model [25, 37]. These defense methods offer varying levels of effectiveness, and the backdoor defense problem remains an active area of research.
3 Degradation of Proxy-based Selection
This section aims to highlight the limitations of sample selection methods that rely on proxy attacks, motivating the development of a proxy-free approach for selecting poisoned samples in backdoor attacks. To provide context, we will first review the general pipeline of backdoor attacks and the use of proxy attack-based sample selection methods
Poisoning-based Backdoor Attacks. Poisoning-based backdoor attacks involve injecting a backdoor into a subset of a clean training set , resulting in a corresponding poisoned set . Here, is a pre-defined poison generator that adds the trigger into a clean sample , while and represent the true label and attack-target label of the clean sample and the poisoning sample , respectively. After poisoning, the victim will train a deep model:
In the above equation, represents the DNN model, and is the corresponding loss function. The poisoned rate can be calculated as . The goal of sample selection methods is to enhance the efficiency of backdoor attacks by selecting an optimal poisoning set . This approach is orthogonal to the mainstream method of improving poisoning efficiency and has not been thoroughly explored yet.
Sample Selection Based on Proxy Attack. Recent studies [39, 43] have highlighted the varying contributions of different samples to backdoor injection, with certain subsets of efficient data exhibiting higher attack success rates. For instance, [39] suggests that the efficiency of a poisoned samples is related to its forgetting events during backdoor injection. To identify the optimal subset, they formulate sample selection as an optimization problem and propose a Filtering-and-Updating Strategy (FUS). Although FUS achieves remarkable performance, it utilizes a proxy attack processing to identify the efficient subset.
As illustrated in Figure 1, poisoning-based backdoor attacks consist of two phases: The attacker combines data (D) and trigger (T) to generate poisoned data (D+T) and releases it to victims, who then use it to train infected DNN models. During training, various data transforms/augmentations (F) and hyperparameters (H)***In this case, H includes any training settings such as the model architecture, the optimizer, and the training hyperparameters. are applied. Typically, the attacker has no access to any information about the victim phase. However, proxy-based sample selection methods rely on the complete attack process to select samples. We argue that all components involved in backdoor attacks, including D, T, F, and H, can affect the efficiency of poisoned samples. When discrepancies arise (e.g., between the transforms and hyperparameters used by the attacker and those used by victims), the effectiveness of the selected samples may be compromised. Therefore, we will first investigate the impact of such discrepancies on proxy-based sample selection methods.
3.1 Experimental Settings
To investigate the proposed question in this section, we conduct experiments using the VGG-16 model [34] as the victim model on the CIFAR-10 dataset [18]. We use the SGD optimizer with an initial learning rate of 0.01, momentum of 0.9, and weight decay of 5e-4, and apply two common data transformations: random crop and random horizontal flip. The training is performed for 70 epochs with a batch size of 256. Poisoned samples are generated using a blending strategy [4] in which the clean samples are blended with a trigger image using the equation , where is set to 0.15. We set the attack target to category 0. We adopt the FUS method with a poisoned rate (the size of the poisoning set is ) and the number of selection iterations following [39]. In accordance with previous literature [22], the attack success rate is defined as the probability of classifying a poisoned test data to the target label. Unless stated otherwise, all experiments in this section follow the aforementioned settings.
3.2 Empirical Studies
Different Data Transforms (F) will decrease the effectiveness of proxy-based sample selection method. We investigated how different data transforms (F) affect the effectiveness of proxy-based sample selection methods. To conduct our analysis, we used the experimental settings described previously and followed them exactly. Specifically, we set the data transform for the proxy poisoning attack in the sample selection phase (FUS) to ”RandomCrop” and ”RandomHorizontalFlip”, and we used ”None”, ”RandomCrop”, ”RandomHorizontalFlip”, ”RandomRotation”, and ”ColorJitter” for the data transform during the actual poisoning attack in the victim phase. Our results, depicted in Figure 2, indicate that different data transforms (F) decrease the effectiveness of proxy-based sample selection methods.
Different Hyperparameter (H) will decrease the effectiveness of proxy-based sample selection method. Furthermore, we explore the impact of different hyperparameter to efficient data searched by FUS [39]. Following the same experimental settings as described above, we conduct experiments using the hyperparameter ”V-16, SGD, 0.01” to search for efficient sets of backdoor attacks and evaluate the attack success rate on several other hyperparameters, including ”V-13, SGD, 0.01”, ”V-13, Adam, 0.001”, ”V-16, Adam, 0.001”, ”R-18, SGD, 0.01”, ”R-18, Adam, 0.001”, and ”P-18, SGD, 0.01”†††V-13, V-16, R-18, and P-18 denote VGG13, VGG-16, ResNet-18, and PreActResNet-18, respectively. . Our results, shown in Figure 3, demonstrate that different hyperparameters (H) can also significantly decrease the effectiveness of the proxy-based sample selection method.
Overall, our results suggest that the efficacy of proxy-based sample selection methods is sensitive to attacker-agnostic data transforms and hyperparameters used during the victim phase. This highlights the importance of using proxy-free search methods to achieve efficient sample selection in backdoor attacks.
Input:
Clean training set ; Size of clean training set ; Backdoor trigger ; Attack target ; Poisoned rate ; Diversity rate ; Pre-trained feature extractor ; Trigger function
Output:
Build the poisoned set ;
4 Methodology
To improve poisoning efficiency, it is important to develop proxy-free sample selection methods that leverage the forensic features of efficient data in poisoning-based backdoor attacks. Before introducing our method, we will first discuss some key observations.
4.1 Forensic Features of Efficient Data in Poisoning-based Backdoor Attacks
We contend that the primary reason for the efficiency of backdoor injection lies in the similarity between clean and corresponding poisoned samples. Typically, the labels of poisoned samples are different from those of the original clean samples, meaning that clean and corresponding poisoned samples share similar features but have completely different labels. In this case, samples with high similarity can be viewed as challenging samples in the poisoning task, making them more efficient than low-similarity samples for backdoor attacks. To validate this hypothesis, we conducted empirical studies on the CIFAR-10 dataset using the same experimental settings as in Sec. 3.1. The left part of Figure 4 depicts the similarity distribution of 500 randomly selected samples and 500 FUS-searched samples [39], which demonstrates that FUS-searched efficient data has higher similarity than random samples. We further examined the contribution of similarity to the data efficiency of backdoor attacks by selecting poisoned samples based on similarity. The right part of Figure 4 shows the attack success rate when different poisoned samples were used. As expected, poisoning with high-similarity samples (Top-500 similarity sampling) resulted in a significantly higher attack success rate than poisoning with low-similarity samples (Bottom-500 similarity sampling). Furthermore, we compared random sampling with similarity-selected sampling. However, both poisoning with high-similarity samples (Top-500 similarity sampling) and low-similarity samples (Bottom-500 similarity sampling) yielded lower attack success rates than poisoning with random samples (Random sampling). This suggests that similarity can only be used to filter out inefficient samples and has no positive correlation with data efficiency in poisoned attacks. We argue that the degradation of Top-500 similarity samples can be attributed to the overly limited diversity of the poisoned samples. The middle part of Figure 4 shows the t-SNE of the 50K whole dataset and the 500 selected samples, indicating that Top-500 similarity samples have limited diversity. Selecting 500 samples randomly from the 5K samples with the highest similarity (Top-5K similarity random 500 sampling) can effectively mitigate this limitation, as shown in the right part of Figure 4. Among them, Top-5K similarity random 500 sampling outperforms random sampling and achieves competitive results with only a slight cost compared to FUS [39] methods. In summary, the similarity between clean and corresponding poisoned samples and the diversity of poisoned sample set are two key factors in the efficient data of backdoor attacks.
4.2 Threat Model
Attacker’s Capacities. Following the basic requirement of poisoning-based backdoor attacks [22, 39], we assume that attackers are permitted to poison a portion of the training data, but are not privy to any other training components in the victim phase, such as the training loss, schedule, or model structure. These assumptions align with potential real-world scenarios in which such threats can occur.
Attacker’s Goals. Our aim is to enhance the efficiency of poisoning attacks without incurring additional costs, by selecting an appropriate poisoning set from the dataset . This approach is distinct from the current methods that enhance efficiency by designing triggers, and hence, our method can be easily incorporated into other attack technologies at minimal cost.
4.3 Proxy-Free Selection Strategy
Motivated by the observation of forensic features that contribute to efficient data in poisoning-based backdoor attacks, as demonstrated in Section 4.1, we propose a simple yet effective sample selection strategy (PFS) for enhancing poisoning efficiency. Unlike previous studies that rely on a proxy backdoor attack for sample selection, our method is based on the similarity between clean and corresponding poisoned samples, which is only dependent on the Data (D) and Trigger (T), and independent of Data Transform (F) and Hyperparameters (H) (proxy-free). Specifically, we utilize a pre-trained feature extractor to compute the cosine similarity (cos()) in the feature space between clean and corresponding poisoned samples. We then select the top most similar samples and randomly sample from this set as the poisoning set, where is a hyper-parameter that controls the diversity of the poisoned samples. The algorithmic procedure of our method is presented in Algorithm 1.
Moreover, we demonstrate that our method can filter out most of the worthless samples, which also enhances the FUS search [39]. Therefore, we propose a combined approach (FUS+PFS) that utilizes the FUS search on the coarse poisoned set containing samples to further enhance the poisoning efficiency.
5 Experiments
We evaluate the effectiveness of PFS on datasets: CIFAR -10 (Sec. 5.2), CIFAR-100 (Sec. 5.3), and Tiny-ImageNet (Sec. A.4 of the Supplementary Materials). We also provide experimental settings in Sec. 5.1, time-consuming analysis in Sec. 5.4, and ablation study in Sec. 5.5.
5.1 Experimental Settings
We conduct experiments with three trigger types (BadNets [13], Blended [4], Optimized [48]), six data transforms (None, RandomCrop, RandomHorizontalFlip, RandomRotation, ColorJitter, RandomCrop+RandomHorizontalFlip), and three DNN architectures (VGG16 [34], ResNet18 [15], PreAct-ResNet18 [16]). We also use four optimizers: SGD-0.01‡‡‡SGD-0.01 means adopting the SGD optimizer with an initial learning rate of 0.01,, SGD-0.02, Adam-0.001, Adam-0.002. The adopted feature extractor is the ResNet-18 that has been pre-trained on the current dataset. Additionally, in Section A.5 of the Supplementary Materials, we present experiments with different pre-trained feature extractors on the CIFAR-10 dataset. Our results show that even using a common feature extractor pre-trained on the ImageNet dataset yields significant improvements compared to the baseline.
We compare our method to the commonly used random selection (Random) [13, 4] and FUS [39]. In the FUS search process, we use VGG16, RandomCrop+RandomHorizontalFlip, and SGD-0.01 to construct the proxy attack. For our method, we use pre-trained ResNet models for similarity measurement and set the diversity hyperparameter to 10 for all three datasets. Other hyperparameters are kept the same as in Sec. 3.1.
| Trigger | Data Transform | Model | |||||||||||
| VGG16 | ResNet18 | PreAct-ResNet18 | |||||||||||
| Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | ||
| BadNets | None | 0.950 | 0.984 | 0.978 | 0.988 | 0.964 | 0.992 | 0.988 | 0.994 | 0.961 | 0.991 | 0.985 | 0.994 |
| RandomCrop | 0.906 | 0.964 | 0.954 | 0.972 | 0.918 | 0.976 | 0.966 | 0.976 | 0.927 | 0.977 | 0.968 | 0.981 | |
| RandomHorizontalFlip | 0.932 | 0.981 | 0.971 | 0.986 | 0.964 | 0.989 | 0.983 | 0.992 | 0.954 | 0.987 | 0.981 | 0.990 | |
| RandomRotation | 0.879 | 0.952 | 0.939 | 0.953 | 0.909 | 0.972 | 0.955 | 0.976 | 0.918 | 0.974 | 0.961 | 0.977 | |
| ColorJitter | 0.952 | 0.982 | 0.975 | 0.987 | 0.967 | 0.993 | 0.988 | 0.995 | 0.964 | 0.992 | 0.985 | 0.994 | |
| RandomCrop+ RandomHorizontalFlip | 0.916 | 0.966 | 0.956 | 0.972 | 0.933 | 0.980 | 0.970 | 0.983 | 0.931 | 0.980 | 0.970 | 0.983 | |
| Avg | 0.923 | 0.972 | 0.962 | 0.976 | 0.943 | 0.984 | 0.975 | 0.986 | 0.943 | 0.984 | 0.975 | 0.987 | |
| Blended | None | 0.795 | 0.856 | 0.861 | 0.889 | 0.814 | 0.923 | 0.885 | 0.930 | 0.867 | 0.949 | 0.919 | 0.957 |
| RandomCrop | 0.792 | 0.830 | 0.865 | 0.874 | 0.818 | 0.888 | 0.885 | 0.915 | 0.821 | 0.903 | 0.896 | 0.921 | |
| RandomHorizontalFlip | 0.827 | 0.899 | 0.916 | 0.930 | 0.867 | 0.948 | 0.925 | 0.959 | 0.864 | 0.954 | 0.923 | 0.962 | |
| RandomRotation | 0.791 | 0.821 | 0.855 | 0.864 | 0.833 | 0.897 | 0.890 | 0.926 | 0.838 | 0.918 | 0.902 | 0.933 | |
| ColorJitter | 0.758 | 0.846 | 0.826 | 0.861 | 0.758 | 0.846 | 0.826 | 0.921 | 0.758 | 0.846 | 0.826 | 0.950 | |
| RandomCrop+ RandomHorizontalFlip | 0.827 | 0.883 | 0.911 | 0.933 | 0.855 | 0.931 | 0.932 | 0.958 | 0.864 | 0.926 | 0.924 | 0.956 | |
| Avg | 0.798 | 0.856 | 0.872 | 0.892 | 0.824 | 0.906 | 0.891 | 0.935 | 0.835 | 0.916 | 0.898 | 0.947 | |
| Optimized | None | 0.844 | 0.973 | 0.910 | 0.988 | 0.840 | 0.952 | 0.897 | 0.968 | 0.902 | 0.982 | 0.940 | 0.993 |
| RandomCrop | 0.952 | 0.998 | 0.990 | 1.000 | 0.966 | 0.999 | 0.994 | 1.000 | 0.970 | 0.999 | 0.993 | 1.000 | |
| RandomHorizontalFlip | 0.922 | 0.990 | 0.975 | 0.998 | 0.885 | 0.977 | 0.945 | 0.993 | 0.932 | 0.994 | 0.981 | 0.998 | |
| RandomRotation | 0.844 | 0.893 | 0.873 | 0.874 | 0.828 | 0.930 | 0.879 | 0.914 | 0.863 | 0.958 | 0.918 | 0.956 | |
| ColorJitter | 0.875 | 0.962 | 0.933 | 0.991 | 0.861 | 0.940 | 0.918 | 0.977 | 0.911 | 0.985 | 0.969 | 0.995 | |
| RandomCrop+ RandomHorizontalFlip | 0.979 | 0.999 | 0.996 | 1.000 | 0.979 | 0.999 | 0.997 | 1.000 | 0.982 | 0.999 | 0.997 | 1.000 | |
| Avg | 0.903 | 0.969 | 0.946 | 0.975 | 0.893 | 0.966 | 0.938 | 0.975 | 0.927 | 0.986 | 0.966 | 0.990 | |
5.2 Experiments on CIFAR-10 Dataset
Results on different data transforms and architectures. Table 1 presents the attack success rate for BadNets, Blended, and Optimized triggers using different data transforms and architectures. We use VGG16, RandomCrop+RandomHorizontalFlip, and SGD-0.01 to construct the proxy attack. Our PFS method significantly improves the efficiency of poisoned data for backdoor attacks. In all 54 settings, the attack success rate of the poisoned samples selected by our method is higher than that of randomly selected samples, at the same poisoning ratios. On average, our method outperforms random selection by 0.044, 0.074, and 0.066 for BadNets, Blended, and Optimized triggers, respectively. In general, our method performs better than FUS, achieving stronger attack strength in 48 out of 54 settings. It is worth noting that the FUS search requires hundreds or even thousands of times more time than our method. Combining FUS with our method yields the best results in 51 out of 54 settings, as our method filters out many inefficient samples and narrows the FUS search space.
Results on different optimizer and architectures. Table 2 shows the attack success rate for the BadNets, Blended, and Optimized triggers with different optimizers and architectures. We use VGG16, RandomCrop+RandomHorizontalFlip, and SGD-0.01 to construct the proxy attack. The results demonstrate the effectiveness of our PFS method. First, for all 36 settings, the attack success rate achieved by our method is higher than that achieved by random selection, with average improvements of 0.036, 0.0437, and 0.017 for the BadNets, Blended, and Optimized triggers, respectively. Second, our method outperforms FUS in 25 out of 36 cases. When combining our method with FUS, we obtain the best results in 33 out of 36 settings.
In the Supplementary Materials, we have included additional experiments to further evaluate the effectiveness and robustness of our proposed strategy. Specifically, in Sec. A.1 and A.2, we present the attack success rate (ASR) and benign accuracy (BA) under poisoned rates of and on the CIFAR-10 dataset. The results demonstrate that our method remains robust to different poisoned rates and target labels while preserving benign accuracy. Additionally, in Sec. A.3, we present the ASR for an additional target class (class 3) on the CIFAR-10 dataset. Overall, our findings confirm the effectiveness of our proposed method and support our hypothesis that the efficiency of backdoor attacks is determined by sample similarity and diversity.
| Trigger | Optimizer | Model | |||||||||||
| VGG16 | ResNet18 | PreAct-ResNet18 | |||||||||||
| Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | ||
| BadNets | SGD-0.01 | 0.916 | 0.966 | 0.956 | 0.972 | 0.933 | 0.980 | 0.970 | 0.983 | 0.931 | 0.980 | 0.970 | 0.983 |
| SGD-0.02 | 0.918 | 0.971 | 0.962 | 0.976 | 0.938 | 0.983 | 0.975 | 0.986 | 0.943 | 0.984 | 0.976 | 0.987 | |
| Adam-0.0005 | 0.920 | 0.934 | 0.963 | 0.971 | 0.945 | 0.960 | 0.975 | 0.985 | 0.944 | 0.960 | 0.976 | 0.984 | |
| Adam-0.001 | 0.928 | 0.971 | 0.967 | 0.957 | 0.950 | 0.985 | 0.977 | 0.986 | 0.954 | 0.982 | 0.977 | 0.985 | |
| Avg | 0.921 | 0.961 | 0.962 | 0.969 | 0.942 | 0.977 | 0.974 | 0.985 | 0.943 | 0.977 | 0.975 | 0.985 | |
| Blended | SGD-0.01 | 0.827 | 0.883 | 0.911 | 0.933 | 0.855 | 0.931 | 0.932 | 0.958 | 0.864 | 0.926 | 0.924 | 0.956 |
| SGD-0.02 | 0.836 | 0.887 | 0.934 | 0.935 | 0.881 | 0.939 | 0.950 | 0.968 | 0.881 | 0.947 | 0.948 | 0.967 | |
| Adam-0.0005 | 0.873 | 0.890 | 0.924 | 0.908 | 0.897 | 0.952 | 0.953 | 0.974 | 0.893 | 0.952 | 0.940 | 0.966 | |
| Adam-0.001 | 0.886 | 0.910 | 0.937 | 0.909 | 0.900 | 0.953 | 0.951 | 0.970 | 0.896 | 0.949 | 0.943 | 0.962 | |
| Avg | 0.881 | 0.891 | 0.927 | 0.921 | 0.883 | 0.944 | 0.947 | 0.968 | 0.884 | 0.944 | 0.939 | 0.963 | |
| Optimized | SGD-0.01 | 0.979 | 0.999 | 0.996 | 1.000 | 0.979 | 0.999 | 0.997 | 1.000 | 0.982 | 0.999 | 0.997 | 1.000 |
| SGD-0.02 | 0.979 | 0.999 | 0.998 | 1.000 | 0.982 | 0.999 | 0.997 | 1.000 | 0.984 | 0.999 | 0.997 | 1.000 | |
| Adam-0.0005 | 0.975 | 0.997 | 0.996 | 1.000 | 0.986 | 0.998 | 0.996 | 1.000 | 0.982 | 0.998 | 0.996 | 1.000 | |
| Adam-0.001 | 0.980 | 0.999 | 0.997 | 1.000 | 0.985 | 0.999 | 0.997 | 1.000 | 0.982 | 0.998 | 0.994 | 1.000 | |
| Avg | 0.978 | 0.999 | 0.997 | 1.000 | 0.983 | 0.999 | 0.997 | 1.000 | 0.983 | 0.997 | 0.996 | 1.000 | |
| Trigger | Data Transform | Model | |||||||||||
| VGG16 | ResNet18 | PreAct-ResNet18 | |||||||||||
| Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | ||
| BadNets | None | 0.933 | 0.970 | 0.965 | 0.971 | 0.950 | 0.978 | 0.973 | 0.979 | 0.949 | 0.977 | 0.970 | 0.977 |
| RandomCrop | 0.882 | 0.947 | 0.930 | 0.945 | 0.894 | 0.951 | 0.940 | 0.952 | 0.888 | 0.951 | 0.940 | 0.951 | |
| RandomHorizontalFlip | 0.920 | 0.964 | 0.958 | 0.965 | 0.936 | 0.970 | 0.965 | 0.971 | 0.929 | 0.968 | 0.962 | 0.968 | |
| RandomRotation | 0.868 | 0.923 | 0.912 | 0.924 | 0.866 | 0.939 | 0.925 | 0.938 | 0.875 | 0.935 | 0.926 | 0.938 | |
| ColorJitter | 0.934 | 0.970 | 0.965 | 0.971 | 0.952 | 0.978 | 0.973 | 0.979 | 0.946 | 0.977 | 0.970 | 0.978 | |
| RandomCrop+ RandomHorizontalFlip | 0.879 | 0.939 | 0.929 | 0.939 | 0.888 | 0.947 | 0.942 | 0.950 | 0.894 | 0.949 | 0.941 | 0.951 | |
| Avg | 0.903 | 0.952 | 0.943 | 0.953 | 0.914 | 0.961 | 0.953 | 0.962 | 0.914 | 0.960 | 0.952 | 0.961 | |
| Blended | None | 0.866 | 0.923 | 0.920 | 0.941 | 0.884 | 0.963 | 0.946 | 0.974 | 0.894 | 0.975 | 0.950 | 0.982 |
| RandomCrop | 0.863 | 0.874 | 0.891 | 0.880 | 0.826 | 0.916 | 0.906 | 0.938 | 0.829 | 0.931 | 0.912 | 0.942 | |
| RandomHorizontalFlip | 0.891 | 0.945 | 0.935 | 0.960 | 0.902 | 0.968 | 0.951 | 0.977 | 0.912 | 0.974 | 0.958 | 0.982 | |
| RandomRotation | 0.818 | 0.862 | 0.901 | 0.865 | 0.841 | 0.916 | 0.916 | 0.936 | 0.848 | 0.926 | 0.921 | 0.944 | |
| ColorJitter | 0.863 | 0.920 | 0.917 | 0.931 | 0.904 | 0.972 | 0.956 | 0.977 | 0.913 | 0.979 | 0.959 | 0.983 | |
| RandomCrop+ RandomHorizontalFlip | 0.811 | 0.865 | 0.905 | 0.872 | 0.826 | 0.918 | 0.920 | 0.942 | 0.850 | 0.931 | 0.924 | 0.947 | |
| Avg | 0.852 | 0.898 | 0.912 | 0.908 | 0.864 | 0.942 | 0.933 | 0.957 | 0.874 | 0.953 | 0.937 | 0.963 | |
| Optimized | None | 0.979 | 0.999 | 0.995 | 0.999 | 0.884 | 0.972 | 0.931 | 0.982 | 0.895 | 0.983 | 0.962 | 0.987 |
| RandomCrop | 0.989 | 0.999 | 0.997 | 1.000 | 0.983 | 0.999 | 0.996 | 1.000 | 0.981 | 0.999 | 0.995 | 1.000 | |
| RandomHorizontalFlip | 0.985 | 0.999 | 0.997 | 1.000 | 0.938 | 0.998 | 0.980 | 0.983 | 0.938 | 0.997 | 0.984 | 0.998 | |
| RandomRotation | 0.810 | 0.831 | 0.867 | 0.785 | 0.827 | 0.909 | 0.882 | 0.882 | 0.835 | 0.908 | 0.883 | 0.900 | |
| ColorJitter | 0.984 | 0.997 | 0.993 | 1.000 | 0.967 | 0.989 | 0.979 | 0.985 | 0.956 | 0.993 | 0.990 | 0.994 | |
| RandomCrop+ RandomHorizontalFlip | 0.991 | 0.999 | 0.997 | 1.000 | 0.987 | 0.999 | 0.998 | 1.000 | 0.982 | 0.998 | 0.994 | 0.999 | |
| Avg | 0.956 | 0.971 | 0.974 | 0.964 | 0.931 | 0.978 | 0.961 | 0.972 | 0.931 | 0.980 | 0.968 | 0.980 | |
| PFS (w/o pre-training) | PFS (w pre-training) | FUS () |
| 17 | 562 | 9,150 |
5.3 Experiments on CIFAR-100 Dataset
The results for CIFAR-100 in Table 3 are similar to those for CIFAR-10. Our proxy-free approach outperforms random selection in all cases, with average boosts of 0.047, 0.068, and 0.037 for BadNets, Blended, and Optimized triggers, respectively. Compared to FUS, PFS achieves better attack results in 49 out of 54 settings.
5.4 Analysis of Time Consumption
Table 4 presents a comparison of the running times of our PFS with those of FUS. Our method demonstrates remarkable speed, being faster than FUS. Even with pre-training time included, our method is still approximately faster. Importantly, our method utilizes the same pre-trained model for different triggers, while FUS needs to conduct a complete search from the beginning for each trigger. This advantage makes our method a practical choice for situations where efficiency is crucial.
5.5 Ablation Study
We conduct an ablation study on the hyperparameter , which determines the diversity of the poisoned sample set. The results are shown in Figure 5, and the attack success rates demonstrate similar trends under different data transforms. Initially, as diversity gradually increases, the attack success rates also increase. However, too large of a value for can reduce the role of similarity and weaken the attack. These findings highlight the importance of balancing similarity and diversity for efficient poisoned samples, supporting our claim in Sec. 4.1.
6 Conclusion and Future Work
Conclusion. This paper presents empirical evidence and reveals two intriguing phenomena related to data efficiency in backdoor attacks. First, we demonstrate that sample selection methods based on proxy attacks suffer from performance degradation when gaps exist between proxy poisoning attacks and actual victim poisoning processes. Second, we show that the similarity between clean and corresponding poisoned samples, as well as the diversity of the poisoned sample set, are key features for efficient data in backdoor attacks. Based on these findings, we propose a simple yet effective proxy-free sample selection strategy that selects poisoned samples based on the similarity between clean and corresponding poisoned samples. Our method significantly improves the poisoning efficiency on benchmark datasets with almost no additional cost.
Future Work. While the proxy-free sampling strategy we proposed has demonstrated efficient data selection in backdoor attacks, we note that the similarity between clean and corresponding poisoned samples does not necessarily guarantee high efficiency, as similarity only filters out non-efficient samples. In the future, it would be desirable to identify a unified indicator that positively correlates with data efficiency in backdoor attacks.
Appendix A Supplementary Materials
This supplementary material is organized as follows. We first discuss the attack success rate and benign accuracy under poisoned rates with and on the CIFAR-10 dataset in Sec. A.1 and Sec. A.2, respectively. Additionally, we presents the results of attack success rate under additional target classes on the CIFAR-10 dataset in Sec. A.3. Besides, Sec. A.4 presents the experimental results on the Tiny-ImageNet dataset. At last, Sec. A.5 shows the experiments on different pre-trained feature extractor .
A.1 Attack Success Rate Under Additional Poisoned Rates on the CIFAR-10 Dataset
We evaluate the attack success rate (ASR) under different poisoned rates on the CIFAR-10 dataset. Tables 5 and 6 present the ASR with trigger BadNets and Blended at poisoned rates of and , respectively. Results under poisoned rates of and are similar to those at .
(i) Experiments under poisoned rates of : Our proxy-free approach outperforms random selection in all cases, with average boosts of 0.030 and 0.055 on BadNets and Blended triggers, respectively. The proxy-free method also achieves better attack results than FUS in most cases (27/36). The combination of these two methods achieves superior results in all 36 settings.
(ii) Experiments under poisoned rates of : Our proxy-free approach outperforms random selection in all cases, with average boosts of 0.023 and 0.04 on BadNets and Blended triggers, respectively. The proxy-free method also achieves better attack results than FUS in most cases (26/36). The combination of these two methods achieves superior results in all 36 settings.
| Trigger | Data Transform | Model | |||||||||||
| VGG16 | ResNet18 | PreAct-ResNet18 | |||||||||||
| Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | ||
| BadNets | None | 0.961 | 0.986 | 0.982 | 0.992 | 0.976 | 0.993 | 0.990 | 0.997 | 0.973 | 0.992 | 0.989 | 0.997 |
| RandomCrop | 0.924 | 0.970 | 0.966 | 0.983 | 0.944 | 0.980 | 0.974 | 0.988 | 0.946 | 0.979 | 0.977 | 0.989 | |
| RandomHorizontalFlip | 0.952 | 0.983 | 0.979 | 0.991 | 0.973 | 0.975 | 0.986 | 0.995 | 0.967 | 0.988 | 0.984 | 0.994 | |
| RandomRotation | 0.915 | 0.968 | 0.958 | 0.977 | 0.929 | 0.979 | 0.971 | 0.988 | 0.941 | 0.979 | 0.973 | 0.988 | |
| ColorJitter | 0.958 | 0.986 | 0.981 | 0.992 | 0.977 | 0.994 | 0.990 | 0.997 | 0.972 | 0.993 | 0.988 | 0.996 | |
| RandomCrop+ RandomHorizontalFlip | 0.932 | 0.970 | 0.965 | 0.981 | 0.949 | 0.980 | 0.978 | 0.989 | 0.952 | 0.981 | 0.978 | 0.989 | |
| Avg | 0.940 | 0.977 | 0.972 | 0.986 | 0.958 | 0.984 | 0.982 | 0.992 | 0.959 | 0.985 | 0.982 | 0.992 | |
| Blended | None | 0.866 | 0.921 | 0.918 | 0.953 | 0.883 | 0.954 | 0.939 | 0.971 | 0.909 | 0.970 | 0.954 | 0.982 |
| RandomCrop | 0.855 | 0.900 | 0.933 | 0.942 | 0.879 | 0.935 | 0.943 | 0.963 | 0.882 | 0.948 | 0.938 | 0.967 | |
| RandomHorizontalFlip | 0.889 | 0.937 | 0.952 | 0.970 | 0.906 | 0.966 | 0.960 | 0.983 | 0.927 | 0.975 | 0.961 | 0.987 | |
| RandomRotation | 0.859 | 0.909 | 0.929 | 0.937 | 0.894 | 0.938 | 0.947 | 0.967 | 0.901 | 0.955 | 0.953 | 0.972 | |
| ColorJitter | 0.851 | 0.915 | 0.912 | 0.938 | 0.872 | 0.950 | 0.929 | 0.969 | 0.913 | 0.966 | 0.945 | 0.976 | |
| RandomCrop+ RandomHorizontalFlip | 0.885 | 0.931 | 0.952 | 0.965 | 0.905 | 0.958 | 0.964 | 0.982 | 0.919 | 0.967 | 0.967 | 0.984 | |
| Avg | 0.868 | 0.919 | 0.933 | 0.951 | 0.890 | 0.950 | 0.947 | 0.973 | 0.909 | 0.964 | 0.953 | 0.978 | |
| Trigger | Data Transform | Model | |||||||||||
| VGG16 | ResNet18 | PreAct-ResNet18 | |||||||||||
| Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | ||
| BadNets | None | 0.971 | 0.987 | 0.986 | 0.993 | 0.982 | 0.994 | 0.992 | 0.998 | 0.977 | 0.993 | 0.991 | 0.997 |
| RandomCrop | 0.940 | 0.973 | 0.970 | 0.986 | 0.954 | 0.980 | 0.979 | 0.991 | 0.946 | 0.982 | 0.979 | 0.992 | |
| RandomHorizontalFlip | 0.961 | 0.984 | 0.981 | 0.992 | 0.976 | 0.991 | 0.988 | 0.995 | 0.975 | 0.989 | 0.986 | 0.995 | |
| RandomRotation | 0.931 | 0.969 | 0.966 | 0.986 | 0.948 | 0.977 | 0.977 | 0.992 | 0.946 | 0.981 | 0.979 | 0.991 | |
| ColorJitter | 0.966 | 0.986 | 0.984 | 0.993 | 0.982 | 0.993 | 0.992 | 0.998 | 0.981 | 0.992 | 0.990 | 0.997 | |
| RandomCrop+ RandomHorizontalFlip | 0.938 | 0.974 | 0.972 | 0.986 | 0.957 | 0.982 | 0.980 | 0.992 | 0.959 | 0.984 | 0.981 | 0.991 | |
| Avg | 0.951 | 0.979 | 0.977 | 0.989 | 0.967 | 0.986 | 0.985 | 0.994 | 0.964 | 0.987 | 0.984 | 0.994 | |
| Blended | None | 0.914 | 0.943 | 0.952 | 0.952 | 0.916 | 0.969 | 0.955 | 0.981 | 0.937 | 0.978 | 0.967 | 0.988 |
| RandomCrop | 0.896 | 0.937 | 0.958 | 0.966 | 0.912 | 0.961 | 0.961 | 0.978 | 0.913 | 0.960 | 0.965 | 0.985 | |
| RandomHorizontalFlip | 0.923 | 0.955 | 0.972 | 0.983 | 0.931 | 0.976 | 0.973 | 0.990 | 0.947 | 0.982 | 0.976 | 0.991 | |
| RandomRotation | 0.900 | 0.936 | 0.952 | 0.963 | 0.915 | 0.958 | 0.958 | 0.973 | 0.922 | 0.965 | 0.965 | 0.982 | |
| ColorJitter | 0.904 | 0.940 | 0.951 | 0.963 | 0.918 | 0.965 | 0.949 | 0.979 | 0.938 | 0.973 | 0.964 | 0.986 | |
| RandomCrop+ RandomHorizontalFlip | 0.922 | 0.955 | 0.973 | 0.986 | 0.927 | 0.968 | 0.977 | 0.991 | 0.933 | 0.973 | 0.979 | 0.992 | |
| Avg | 0.910 | 0.944 | 0.960 | 0.969 | 0.920 | 0.966 | 0.962 | 0.969 | 0.932 | 0.972 | 0.969 | 0.987 | |
A.2 Benign Accuracy Under Additional Poisoned Rates on the CIFAR-10 Dataset
We also present the benign accuracy (BA) (i.e., the probability of classifying a benign test data to the correct label) under other poisoned rates on the CIFAR-10 dataset. Table 7 and Table 8 illustrate the benign accuracy with trigger BadNets and Blended under poisoned rates of and , respectively. The results show the similar BA with four strategies, which demonstrates that our proposed strategy has no negative impact on benign accuracy.
| Trigger | Data Transform | Model | |||||||||||
| VGG16 | ResNet18 | PreAct-ResNet18 | |||||||||||
| Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | ||
| BadNets | None | 0.868 | 0.868 | 0.865 | 0.866 | 0.837 | 0.837 | 0.838 | 0.835 | 0.850 | 0.849 | 0.848 | 0.847 |
| RandomCrop | 0.903 | 0.905 | 0.905 | 0.903 | 0.914 | 0.912 | 0.914 | 0.911 | 0.914 | 0.914 | 0.914 | 0.914 | |
| RandomHorizontalFlip | 0.892 | 0.892 | 0.893 | 0.894 | 0.878 | 0.879 | 0.879 | 0.878 | 0.886 | 0.888 | 0.886 | 0.883 | |
| RandomRotation | 0.876 | 0.873 | 0.876 | 0.874 | 0.883 | 0.881 | 0.880 | 0.880 | 0.889 | 0.888 | 0.889 | 0.889 | |
| ColorJitter | 0.860 | 0.860 | 0.860 | 0.859 | 0.833 | 0.829 | 0.833 | 0.830 | 0.843 | 0.844 | 0.841 | 0.843 | |
| RandomCrop+ RandomHorizontalFlip | 0.920 | 0.919 | 0.920 | 0.921 | 0.927 | 0.926 | 0.928 | 0.925 | 0.927 | 0.928 | 0.929 | 0.927 | |
| Avg | 0.887 | 0.886 | 0.887 | 0.886 | 0.879 | 0.877 | 0.879 | 0.877 | 0.885 | 0.885 | 0.885 | 0.884 | |
| Blended | None | 0.865 | 0.866 | 0.865 | 0.866 | 0.836 | 0.832 | 0.832 | 0.833 | 0.846 | 0.846 | 0.849 | 0.847 |
| RandomCrop | 0.904 | 0.906 | 0.904 | 0.904 | 0.912 | 0.913 | 0.912 | 0.913 | 0.913 | 0.914 | 0.915 | 0.914 | |
| RandomHorizontalFlip | 0.892 | 0.892 | 0.889 | 0.891 | 0.876 | 0.878 | 0.875 | 0.875 | 0.882 | 0.884 | 0.885 | 0.887 | |
| RandomRotation | 0.873 | 0.875 | 0.875 | 0.873 | 0.880 | 0.880 | 0.881 | 0.880 | 0.887 | 0.888 | 0.889 | 0.889 | |
| ColorJitter | 0.858 | 0.859 | 0.860 | 0.858 | 0.829 | 0.832 | 0.824 | 0.828 | 0.839 | 0.841 | 0.843 | 0.842 | |
| RandomCrop+ RandomHorizontalFlip | 0.922 | 0.920 | 0.919 | 0.920 | 0.928 | 0.927 | 0.928 | 0.928 | 0.928 | 0.927 | 0.929 | 0.928 | |
| Avg | 0.886 | 0.886 | 0.885 | 0.885 | 0.877 | 0.877 | 0.875 | 0.876 | 0.883 | 0.883 | 0.885 | 0.885 | |
| Trigger | Data Transform | Model | |||||||||||
| VGG16 | ResNet18 | PreAct-ResNet18 | |||||||||||
| Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | ||
| BadNets | None | 0.867 | 0.865 | 0.866 | 0.863 | 0.837 | 0.835 | 0.835 | 0.833 | 0.851 | 0.846 | 0.848 | 0.847 |
| RandomCrop | 0.904 | 0.905 | 0.905 | 0.905 | 0.913 | 0.914 | 0.912 | 0.912 | 0.916 | 0.915 | 0.913 | 0.913 | |
| RandomHorizontalFlip | 0.891 | 0.891 | 0.891 | 0.891 | 0.878 | 0.878 | 0.878 | 0.876 | 0.886 | 0.885 | 0.886 | 0.885 | |
| RandomRotation | 0.873 | 0.876 | 0.874 | 0.872 | 0.882 | 0.882 | 0.881 | 0.880 | 0.888 | 0.889 | 0.889 | 0.890 | |
| ColorJitter | 0.862 | 0.857 | 0.857 | 0.859 | 0.830 | 0.827 | 0.829 | 0.831 | 0.844 | 0.843 | 0.843 | 0.841 | |
| RandomCrop+ RandomHorizontalFlip | 0.920 | 0.920 | 0.919 | 0.919 | 0.928 | 0.926 | 0.926 | 0.925 | 0.928 | 0.930 | 0.928 | 0.927 | |
| Avg | 0.886 | 0.886 | 0.885 | 0.885 | 0.878 | 0.877 | 0.877 | 0.876 | 0.886 | 0.885 | 0.885 | 0.884 | |
| Blended | None | 0.865 | 0.866 | 0.865 | 0.866 | 0.836 | 0.832 | 0.832 | 0.833 | 0.846 | 0.846 | 0.849 | 0.847 |
| RandomCrop | 0.904 | 0.906 | 0.904 | 0.904 | 0.912 | 0.913 | 0.912 | 0.913 | 0.913 | 0.914 | 0.915 | 0.914 | |
| RandomHorizontalFlip | 0.892 | 0.892 | 0.889 | 0.891 | 0.876 | 0.878 | 0.875 | 0.875 | 0.882 | 0.884 | 0.885 | 0.887 | |
| RandomRotation | 0.873 | 0.875 | 0.875 | 0.873 | 0.880 | 0.880 | 0.881 | 0.880 | 0.887 | 0.888 | 0.889 | 0.889 | |
| ColorJitter | 0.858 | 0.859 | 0.860 | 0.858 | 0.829 | 0.832 | 0.824 | 0.828 | 0.839 | 0.841 | 0.843 | 0.842 | |
| RandomCrop+ RandomHorizontalFlip | 0.922 | 0.920 | 0.919 | 0.920 | 0.928 | 0.927 | 0.928 | 0.928 | 0.928 | 0.927 | 0.929 | 0.928 | |
| Avg | 0.886 | 0.886 | 0.885 | 0.885 | 0.877 | 0.877 | 0.875 | 0.876 | 0.883 | 0.883 | 0.885 | 0.885 | |
A.3 Attack Success Rate Under Additional Target Classes on the CIFAR-10 Dataset
This section presents the attack success rate (ASR) under additional target classes on the CIFAR-10 dataset. Table 9 illustrates the attack success rate with trigger BadNets and Blended under target category 3. Compared with random selection, our proxy-free approach outperforms in all cases, with average boosts of 0.045 and 0.076 on BadNets and Blended triggers. The proxy-free method also achieves better attack results in most cases compared to FUS (22/24). The combination of these two achieves superior results in all 24 settings. These outcomes consistently confirm the conclusions we reached in the paper.
| Trigger | Data Transform | Model | |||||||
| VGG16 | ResNet18 | ||||||||
| Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | ||
| BadNets | None | 0.950 | 0.984 | 0.977 | 0.986 | 0.959 | 0.991 | 0.985 | 0.994 |
| RandomCrop | 0.909 | 0.960 | 0.954 | 0.968 | 0.925 | 0.976 | 0.966 | 0.979 | |
| RandomHorizontalFlip | 0.934 | 0.980 | 0.972 | 0.983 | 0.959 | 0.987 | 0.981 | 0.990 | |
| RandomRotation | 0.890 | 0.951 | 0.942 | 0.951 | 0.912 | 0.968 | 0.955 | 0.972 | |
| ColorJitter | 0.945 | 0.983 | 0.975 | 0.986 | 0.966 | 0.991 | 0.985 | 0.993 | |
| RandomCrop+ RandomHorizontalFlip | 0.905 | 0.964 | 0.957 | 0.968 | 0.925 | 0.978 | 0.968 | 0.980 | |
| Avg | 0.922 | 0.970 | 0.963 | 0.974 | 0.941 | 0.982 | 0.973 | 0.985 | |
| Blended | None | 0.774 | 0.867 | 0.851 | 0.899 | 0.806 | 0.915 | 0.850 | 0.935 |
| RandomCrop | 0.780 | 0.813 | 0.822 | 0.869 | 0.791 | 0.902 | 0.827 | 0.918 | |
| RandomHorizontalFlip | 0.836 | 0.899 | 0.887 | 0.934 | 0.845 | 0.949 | 0.888 | 0.962 | |
| RandomRotation | 0.793 | 0.832 | 0.813 | 0.847 | 0.804 | 0.898 | 0.840 | 0.903 | |
| ColorJitter | 0.778 | 0.836 | 0.805 | 0.874 | 0.814 | 0.907 | 0.854 | 0.927 | |
| RandomCrop+ RandomHorizontalFlip | 0.815 | 0.843 | 0.871 | 0.884 | 0.838 | 0.927 | 0.893 | 0.951 | |
| Avg | 0.796 | 0.848 | 0.842 | 0.885 | 0.816 | 0.916 | 0.859 | 0.933 | |
A.4 Experiments on Tiny-ImageNet Dataset
In Table 10, it can be observed that our method outperforms the random method in terms of attack success rates of poisoned samples on Tiny-ImageNet, with average boosts of 0.054, 0.094, and 0.072 on BadNets, Blended, and Optimized triggers, respectively. In comparison to FUS, PFS outperforms FUS in 45 out of 54 cases. When the two methods are combined, the best results are achieved in 34 out of 54 settings, indicating the superiority of our method. These findings provide strong evidence supporting the effectiveness of our proposed method.
| Trigger | Data Transform | Model | |||||||||||
| VGG16 | ResNet18 | PreAct-ResNet18 | |||||||||||
| Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | ||
| BadNets | None | 0.952 | 0.990 | 0.983 | 0.989 | 0.960 | 0.992 | 0.985 | 0.991 | 0.956 | 0.991 | 0.981 | 0.989 |
| RandomCrop | 0.893 | 0.959 | 0.950 | 0.958 | 0.917 | 0.970 | 0.963 | 0.973 | 0.904 | 0.973 | 0.958 | 0.971 | |
| RandomHorizontalFlip | 0.952 | 0.989 | 0.984 | 0.989 | 0.963 | 0.991 | 0.981 | 0.989 | 0.947 | 0.983 | 0.974 | 0.985 | |
| RandomRotation | 0.881 | 0.932 | 0.937 | 0.944 | 0.912 | 0.959 | 0.964 | 0.966 | 0.910 | 0.957 | 0.963 | 0.970 | |
| ColorJitter | 0.956 | 0.990 | 0.985 | 0.990 | 0.966 | 0.993 | 0.986 | 0.992 | 0.956 | 0.990 | 0.983 | 0.988 | |
| RandomCrop+ RandomHorizontalFlip | 0.900 | 0.943 | 0.955 | 0.950 | 0.917 | 0.977 | 0.967 | 0.975 | 0.924 | 0.974 | 0.965 | 0.978 | |
| Avg | 0.922 | 0.967 | 0.966 | 0.970 | 0.939 | 0.980 | 0.974 | 0.981 | 0.933 | 0.978 | 0.971 | 0.980 | |
| Blended | None | 0.889 | 0.964 | 0.947 | 0.975 | 0.876 | 0.965 | 0.930 | 0.973 | 0.873 | 0.969 | 0.941 | 0.978 |
| RandomCrop | 0.836 | 0.928 | 0.920 | 0.942 | 0.835 | 0.931 | 0.914 | 0.936 | 0.822 | 0.927 | 0.906 | 0.932 | |
| RandomHorizontalFlip | 0.905 | 0.966 | 0.964 | 0.980 | 0.888 | 0.964 | 0.950 | 0.977 | 0.886 | 0.965 | 0.954 | 0.981 | |
| RandomRotation | 0.865 | 0.931 | 0.946 | 0.953 | 0.872 | 0.945 | 0.942 | 0.966 | 0.870 | 0.950 | 0.941 | 0.966 | |
| ColorJitter | 0.883 | 0.958 | 0.945 | 0.964 | 0.871 | 0.956 | 0.917 | 0.962 | 0.872 | 0.967 | 0.931 | 0.971 | |
| RandomCrop+ RandomHorizontalFlip | 0.850 | 0.931 | 0.940 | 0.958 | 0.840 | 0.929 | 0.919 | 0.942 | 0.828 | 0.931 | 0.912 | 0.940 | |
| Avg | 0.871 | 0.946 | 0.944 | 0.962 | 0.864 | 0.948 | 0.929 | 0.959 | 0.859 | 0.952 | 0.931 | 0.961 | |
| Optimized | None | 0.907 | 0.981 | 0.956 | 0.984 | 0.875 | 0.968 | 0.934 | 0.968 | 0.862 | 0.970 | 0.935 | 0.967 |
| RandomCrop | 0.897 | 0.976 | 0.959 | 0.980 | 0.886 | 0.968 | 0.950 | 0.972 | 0.869 | 0.958 | 0.945 | 0.951 | |
| RandomHorizontalFlip | 0.903 | 0.979 | 0.956 | 0.985 | 0.892 | 0.976 | 0.950 | 0.976 | 0.877 | 0.975 | 0.948 | 0.974 | |
| RandomRotation | 0.899 | 0.977 | 0.957 | 0.972 | 0.897 | 0.978 | 0.956 | 0.972 | 0.888 | 0.973 | 0.948 | 0.956 | |
| ColorJitter | 0.695 | 0.714 | 0.744 | 0.707 | 0.769 | 0.773 | 0.803 | 0.775 | 0.772 | 0.779 | 0.816 | 0.781 | |
| RandomCrop+ RandomHorizontalFlip | 0.892 | 0.973 | 0.960 | 0.974 | 0.890 | 0.964 | 0.952 | 0.968 | 0.870 | 0.963 | 0.949 | 0.957 | |
| Avg | 0.866 | 0.933 | 0.922 | 0.934 | 0.868 | 0.938 | 0.924 | 0.939 | 0.856 | 0.936 | 0.924 | 0.931 | |
A.5 Experiments on Different Pre-trained Feature Extractor
To compute the cosine similarity in the proposed PFS, a pre-trained feature extractor is required. This section examines the impact of various pre-trained feature extractors on the CIFAR-10 dataset. As demonstrated in Figure 6, even using a common feature extractor pre-trained on the ImageNet dataset yields significant improvements compared to the baseline.

### v2

@@
-A Proxy-Free Strategy for Practically Improving the Poisoning Efficiency in Backdoor Attacks
+A Proxy Attack-Free Strategy for Practically Improving the Poisoning Efficiency in Backdoor Attacks
 Abstract
-Poisoning efficiency is a crucial factor in poisoning-based backdoor attacks.
-Attackers prefer to use as few poisoned samples as possible to achieve the same level of attack strength, in order to remain undetected.
-Efficient triggers have significantly improved poisoning efficiency, but there is still room for improvement.
-Recently, selecting efficient samples has shown promise, but it requires a proxy backdoor injection task to find an efficient poisoned sample set, which can lead to performance degradation if the proxy attack settings are different from the actual settings used by the victims.
-In this paper, we propose a novel Proxy-Free Strategy (PFS) that selects efficient poisoned samples based on individual similarity and set diversity, effectively addressing this issue.
-We evaluate the proposed strategy on several datasets, triggers, poisoning ratios, architectures, and training hyperparameters.
-Our experimental results demonstrate that PFS achieves higher backdoor attack strength while 500 faster than previous proxy-based selection approaches.
-1 Introduction
+Poisoning efficiency plays a critical role in poisoning-based backdoor attacks.
+To evade detection, attackers aim to use the fewest poisoning samples while achieving the desired attack strength.
+Although efficient triggers have significantly improved poisoning efficiency, there is still room for further enhancement.
+Recently, selecting efficient samples has shown promise, but it often requires a proxy backdoor injection task to identify an efficient poisoning sample set.
+However, the proxy attack-based approach can lead to performance degradation if the proxy attack settings differ from those used by the actual victims due to the shortcut of backdoor learning.
+This paper presents a Proxy attack-Free Strategy (PFS) designed to identify efficient poisoning samples based on individual similarity and ensemble diversity, effectively addressing the mentioned concern.
+The proposed PFS is motivated by the observation that selecting the to-be-poisoned samples with high similarity between clean samples and their corresponding poisoning samples results in significantly higher attack success rates compared to using samples with low similarity.
+Furthermore, theoretical analyses for this phenomenon are provided based on the theory of active learning and neural tangent kernel.
+We comprehensively evaluate the proposed strategy across various datasets, triggers, poisoning rates, architectures, and training hyperparameters.
+Our experimental results demonstrate that PFS enhances backdoor attack efficiency, while also exhibiting a remarkable speed advantage over prior proxy-dependent selection methodologies.
+Index Terms:
+Backdoor attack, Sample selection.I Introduction
 The abundance of training data plays a vital role in the success of Deep Neural Networks (DNNs).
 For instance, GPT-3 [3], a deep learning model with 175 billion parameters, owes its effectiveness in various language processing tasks to its pretraining on a massive corpus of 45 TB of text data.
-With the increasing demand for data, users and businesses are turning to third-party sources or online collection as easier means of obtaining it.
-However, recent studies [4, 12, 3, 11] have demonstrated that such practices can be maliciously exploited by attackers to contaminate training data, thereby negatively impacting the functionality and reliability of trained models.
-During the training phase, one of the significant threats is called backdoor attacks [4, 12, 27].
-This attack involves injecting a covert backdoor into the DNN by introducing a small number of poisoned samples into the benign training set.
-Although the model appears normal on benign inputs, a predefined trigger can activate the infected model, forcing its predictions to align with the attacker’s objective.
+Responding to the increasing demand for data, both users and businesses are increasingly turning to third-party sources or online repositories, which offer more convenient way for data collection.
+However, recent researches [4, 14, 13] have demonstrated that such practices can be maliciously exploited by attackers to poison training data, thereby negatively impacting the functionality and reliability of trained models.
+During the training phase, a significant threat emerges in the form of backdoor attacks [4, 14, 33].
+This attack involves injecting a covert backdoor into the DNNs by introducing a small number of poisoning samples into an otherwise benign training dataset.
+While the model appears normal when presented with benign inputs, a predefined trigger can activate the infected model, forcing its predictions to align with the objective of attackers.
 As research continues to reveal this threat in various tasks such as speaker verification [59] and malware detection [24], attention to backdoor attacks is increasing.
-Poisoning efficiency is a crucial indicator for attackers attempting to launch data poisoning attacks on DNNs [47, 22, 39, 40].
-The aim is to contaminate the training dataset with as few poisoned samples as possible while achieving the desired outcome.
-A perfect scenario is if a backdoor can be successfully implanted by poisoning only one sample in the dataset, significantly enhancing the attack’s stealthiness.
-Currently, methods to improve poisoning efficiency fall into two categories: designing efficient triggers [47, 48, 44] and selecting efficient samples [39].
-The former aims to find trigger patterns that are easier for models to learn, while the latter focuses on selecting and representative poisoned samples.
-This paper is concerned with the latter approach.
-A recent study [39] on sample selection has yielded significant results, demonstrating that the same attack success rate can be achieved with only to of the poisoned samples compared to random selection strategies.
-However, we have found that this method requires a proxy backdoor injection task to evaluate the contribution of each poisoned sample, resulting in a constructed poisoned set closely related to the attacker’s specific proxy attack.
-This presents a challenge, as the effectiveness of the selected set will be damaged when the attacker’s proxy attack settings differ from the actual settings of the victims, which is a common scenario in practice.
-Moreover, conducting the proxy task significantly increases the execution time of the algorithm, limiting its scalability on large datasets.
-In this paper, we address the above issue and present a Proxy-Free Strategy (PFS) that acquires higher backdoor attack strength while 500 faster than previous selection method.
-Our main contributions are as follows:
-We conduct an analysis of the forensic features of efficient poisoned samples in backdoor attacks, with a specific focus on two key factors: the similarity between clean and corresponding poisoned samples, and the diversity of the poisoned sample set.
-Based on this analysis, we propose a novel selection method called the Proxy-Free Strategy (PFS), which does not require a proxy attack task, thereby addressing the issue of inconsistent settings and significantly reducing the time consumption.
-we evaluate PFS on three datasets: CIFAR-10, CIFAR-100, and Tiny-ImageNet, and our results consistently demonstrate its effectiveness.
-Notably, PFS achieves comparable or even higher attack success rates while 500 faster than previous proxy-based selection approaches, making it more scalable for larger datasets or complex models.
-2 Related Works
-2.1 Backdoor Attacks
-Backdoor attacks aim to embed hidden Trojans into DNNs, enabling the attacked models to perform well on clean samples while misbehaving on samples with specifically designed triggers.
-These attacks can occur at any stage of DNN development, including code [1, 33], outsourcing, pretrained [36, 10, 19], data collection [12, 4, 35, 26], collaborative learning [42, 31], and post-deployment [32, 17].
-Among these, poisoning-based backdoor attacks, which involve injecting backdoor triggers by modifying the training data, are the most straightforward and widely adopted.
-Recent studies have focused on improving poisoning efficiency, which can be divided into two categories.
-Designing Efficient Trigger.
-Trigger design is a crucial aspect of backdoor attacks.
+Poisoning efficiency stands as a pivotal metric for attackers engaging in backdoor attacks against DNNs [60, 26, 50, 52].
+Their objective revolves around poisoning the training dataset with minimal poisoning samples while achieving the desired outcomes.
+An ideal scenario is one where a single poisoning sample suffices to implant a backdoor, greatly enhancing the stealthiness of the attack.
+Currently, strategies aimed at enhancing poisoning efficiency fall into two main categories: designing efficient triggers [60, 61, 57] and selecting efficient samples [50].
+The former focuses on identifying trigger patterns that are easier for models to learn, while the latter centers around the careful selection of representative samples for poisoning.
+This paper aligns with the latter approach, exploring the differences in poisoning efficiency of different poisoning samples in backdoor attacks.
+A recent investigation [50] into sample selection has yielded significant results, demonstrating that an equivalent attack success rate can be achieved using only to of the poisoning samples in comparison to random selection strategies.
+However, it has become apparent that this method relies on a proxy backdoor task to assess the contribution of each poisoning sample, thus constructing an efficient poison set closely associated with the specific proxy attack of the attacker.
+This proxy attack-based selection method [50] poses a challenge due to potential deterioration in its effectiveness when the proxy attack settings of attackers deviate from the actual settings of the victims—a frequent attack scenario [60, 61, 57].
+Moreover, conducting the proxy attack task significantly increases the execution time of the algorithm, limiting its scalability on large datasets.
+This paper presents a solution to the aforementioned challenge by introducing a novel Proxy attack-Free Strategy (PFS) that not only achieves higher backdoor attack efficiency but also boasts a remarkable speedup compared to previous selection methods.
+Our contributions can be summarized as follows:
+This paper makes a comprehensive study of the forensic properties inherent in efficient poisoning samples within the context of backdoor attacks.
+In particular, we delve into two critical facets: the similarity between clean samples and their corresponding poisoning samples, and the diversity of poisoning set.
+Drawing on the insights derived from our analysis, we introduce a novel selection approach named the Proxy attack-Free Strategy (PFS).
+Distinguishing itself from previous study, PFS eliminates the requirement for a proxy attack task, effectively addressing the challenge arising from differing settings while significantly reducing the time overhead.
+The effectiveness of PFS is subjected to evaluation across three diverse datasets: CIFAR-10, CIFAR-100, and Tiny-ImageNet.
+Our empirical findings consistently confirm its efficacy.
+It is worth noting that PFS has achieved superior attack success rates while achieving a remarkable acceleration when contrasted with prior proxy-dependent selection methodologies.
+This attribute renders PFS particularly well-suited for scalability, making it apt for accommodating larger datasets and complex models.
+II Related Works
+II-A Backdoor Attacks
+Backdoor attacks aim to inject Trojans within DNNs.
+This manipulation empowers poisoning models to yield accurate outcomes when presented with clean samples, while displaying anomalous behavior with samples containing designed triggers.
+These attacks can occur at various stages within the DNN development lifecycle, including code [1, 41], outsourcing [51], pre-trained models [45, 12], data collection [14, 4, 44, 31], collaborative learning [54, 39], and even post-deployment scenarios [40, 21].
+Of these ways, the most straightforward and widely adopted approach is poisoning-based backdoor attacks, entailing the insertion of backdoor triggers through modifications to the training data.
+Recent researches on poisoning-based backdoor attacks have been proposed to enhancing poisoning efficiency from two aspects.
+Designing Efficient Triggers.
+Trigger design is a hot topic in backdoor learning.
 Chen et al.
-[4] proposed a blended strategy that blends clean samples with the trigger to evade human detection.
-Later, studies aimed to create more efficient and invisible triggers by adopting natural patterns such as warping [30], rotation [38], style transfer [5], frequency [8, 45], and reflection [28].
-In some studies, researchers have taken inspiration from Universal Adversarial Perturbations (UAPs) [29] and optimized an UAP on a pre-trained clean model to create a trigger, which is an effective and widely used approach [48, 21, 6].
-Li et al.
-[23] proposed a different approach where they use GANs models to acquire a sample-specific trigger, which is a departure from previous works that adopt universal triggers.
+[4] initially proposed a blended strategy for evading human detection, acquiring poisoning samples by blending clean samples with triggers.
+Subsequent studies have explored leveraging natural patterns such as warping [38], rotation [48], style transfer [5], frequency [9, 58], and reflection [34] to create triggers that are more imperceptible and efficient compared to previous methods.
+Drawing inspiration from Universal Adversarial Perturbations (UAPs) [37] in adversarial examples, some research [61, 25, 7] optimizes an UAP on a pre-trained clean model as the trigger, a highly effective and widely adopted approach.
+Unlike previous approaches that use universal triggers, Li et al.
+[27] employ GANs models to generate sample-specific triggers.
 Selecting Efficient Poisoning Samples.
-Efficient sample selection for poisoning attacks is a crucial but under-explored area, and is separate from the trigger design.
+Efficient poisoning samples selection is an under-explored area and is orthogonal to the trigger design.
 Xia et al.
-[39] were among the first to explore the contribution of different data to backdoor injection, discovering that not all poisoned samples contribute equally and that appropriate sample selection can significantly improve the efficiency of data in backdoor attacks.
-However, their study [39] relies on proxy attacks to identify the efficient subset, which suffers from performance degradation when gaps exist (e.g., in Transform (F) and Hyperparameters (H) as shown in Figure 1 between the proxy poisoning attack and the actual poisoning process used by victims.
+[50] first explored the contribution to backdoor injection for different data.
+Their work illustrated the inequitable influence of individual poisoning samples and underscored the enormous potential for improved data efficiency within backdoor attacks through efficient sample selection.
+Simultaneously, Gao et al.
+[10] assert a similar notion—that not all samples equally foster the process of poisoning in backdoor attacks.
+They introduce the concept of ”forgetting events” to spotlight the most potent poisoning samples in the context of clean-label backdoor attacks.
+However, both of these studies [50, 10] rely on proxy attacks to determine the efficient subset of samples.
+Regrettably, this dependence on proxy attacks introduces a vulnerability, leading to a degradation in attack performance when disparities arise in parameters like Transformation (F) and Hyperparameters (H) between the proxy poisoning attack and the actual poisoning process, as depicted in Figure 1.
 Therefore, a sample selection strategy that is not based on proxy attacks is urgently needed, and is the main contribution of our paper.
-2.2 Backdoor Defenses
-Backdoor attacks pose a significant threat to the security of DNN models, and several defense methods have been proposed to mitigate these threats.
-Existing backdoor defenses can be broadly classified into two categories: backdoor detection and backdoor erasing.
-Backdoor detection aims to detect poisoning in input data or DNN models used [9, 14, 7, 41].
-While these methods have demonstrated promising detection precision, they cannot completely mitigate the potential threats of backdoor models.
-Backdoor erasing aims to purify the threatened models by erasing the backdoor during or post training.
-During training, clean models can be trained on poisoned data, as proposed by Li et al.
+Comparable to our approach, the Outlier Poisoning Strategy (OPS) [17] also utilizes a surrogate model for efficient sample selection in backdoor attacks against anti-spoof rebroadcast detection.
+The primary concept behind this strategy involves embedding triggers into samples that pose the greatest classification challenge (referred to as outlier samples), encouraging the network to depend on the trigger for classification.
+However, OPS [17] is specifically designed for binary classification and clean-label settings, exhibiting diminished performance in scenarios involving dirty-label settings and multiple classifications.
+II-B Backdoor Defenses
+Backdoor attacks pose a significant threat to the security of Deep Neural Network (DNN), leading to the emergence of various defense mechanisms aimed at mitigating these risks.
+The current landscape of backdoor defenses can be broadly categorized into two classes: backdoor detection and backdoor erasure.
+Backdoor detection aims to identify the presence of poisoning either in input data or within the employed DNN models themselves [11, 18, 8, 53].
+Conversely, backdoor erasure strategies aim to repair poisoning models by removing the embedded backdoor, either during or after the training process.
+In the training phase, the concept of training clean models on poisoned data has been introduced, exemplified by Li et al.
 [28].
-Post-training, models can be fine-tuned, distilled, or pruned to erase the backdoor from the attacked model [25, 37].
-These defense methods offer varying levels of effectiveness, and the backdoor defense problem remains an active area of research.
-3 Degradation of Proxy-based Selection
-This section aims to highlight the limitations of sample selection methods that rely on proxy attacks, motivating the development of a proxy-free approach for selecting poisoned samples in backdoor attacks.
-To provide context, we will first review the general pipeline of backdoor attacks and the use of proxy attack-based sample selection methods
+Post-training, a range of techniques, including fine-tuning, distillation, and pruning, can be employed to eliminate the backdoor from the poisoning model [29, 47].
+III Degradation of proxy attack-based Selection
+The purpose of this section is to emphasize the limitations associated with sample selection methods that rely on proxy attacks, which serves as a motivation for designing a proxy attack-free sample selection strategy in backdoor attacks.
+To better understand our key observations, we will first review the pipeline of backdoor attacks and sample selection methods based on proxy attack.
 Poisoning-based Backdoor Attacks.
 Poisoning-based backdoor attacks involve injecting a backdoor into a subset of a clean training set , resulting in a corresponding poisoned set .
-Here, is a pre-defined poison generator that adds the trigger into a clean sample , while and represent the true label and attack-target label of the clean sample and the poisoning sample , respectively.
+Here, is a pre-defined generator that adds the trigger into clean sample , while and represent the true label and attack-target label of the clean sample and the poisoning sample , respectively.
 After poisoning, the victim will train a deep model:
 In the above equation, represents the DNN model, and is the corresponding loss function.
-The poisoned rate can be calculated as .
+The poisoning rate can be calculated as .
 The goal of sample selection methods is to enhance the efficiency of backdoor attacks by selecting an optimal poisoning set .
 This approach is orthogonal to the mainstream method of improving poisoning efficiency and has not been thoroughly explored yet.
-Sample Selection Based on Proxy Attack.
-Recent studies [39, 43] have highlighted the varying contributions of different samples to backdoor injection, with certain subsets of efficient data exhibiting higher attack success rates.
-For instance, [39] suggests that the efficiency of a poisoned samples is related to its forgetting events during backdoor injection.
-To identify the optimal subset, they formulate sample selection as an optimization problem and propose a Filtering-and-Updating Strategy (FUS).
-Although FUS achieves remarkable performance, it utilizes a proxy attack processing to identify the efficient subset.
-As illustrated in Figure 1, poisoning-based backdoor attacks consist of two phases: The attacker combines data (D) and trigger (T) to generate poisoned data (D+T) and releases it to victims, who then use it to train infected DNN models.
-During training, various data transforms/augmentations (F) and hyperparameters (H)***In this case, H includes any training settings such as the model architecture, the optimizer, and the training hyperparameters.
-are applied.
-Typically, the attacker has no access to any information about the victim phase.
-However, proxy-based sample selection methods rely on the complete attack process to select samples.
-We argue that all components involved in backdoor attacks, including D, T, F, and H, can affect the efficiency of poisoned samples.
-When discrepancies arise (e.g., between the transforms and hyperparameters used by the attacker and those used by victims), the effectiveness of the selected samples may be compromised.
-Therefore, we will first investigate the impact of such discrepancies on proxy-based sample selection methods.
-3.1 Experimental Settings
-To investigate the proposed question in this section, we conduct experiments using the VGG-16 model [34] as the victim model on the CIFAR-10 dataset [18].
-We use the SGD optimizer with an initial learning rate of 0.01, momentum of 0.9, and weight decay of 5e-4, and apply two common data transformations: random crop and random horizontal flip.
+Proxy Attacks-based Sample Selection.
+Recent investigations [50, 55, 10] have highlighted that diverse samples contribute differently to the backdoor injection and a particular subset with efficient data has higher attack success rate.
+Notably, [50, 10] claim that the efficacy of poisoning samples is intertwined with instances of ”forgetting events” during the backdoor injection process.
+To distill the most efficient subset, [50]***It is worth noting that both [50] and [10] employ the concept of forgetting events to guide the selection of poisoned sets.
+It is worth highlighting that while FUS [50] employs multiple iterations of screening, [10] adopts a one-step screening approach, rendering [10] a specific case of FUS.
+As such, for the purpose of analyzing the drawbacks of proxy attack-based methods for selecting poisoning samples, we treat FUS as a representative example in this section.
+formulates the sample selection as an optimization problem and introduce a Filtering-and-Updating Strategy (FUS).
+Although FUS has acquired remarkable performance, it uses a proxy attack processing to find the efficient subset.
+As illustrated in Figure 1, poisoning-based backdoor attacks contain two phases: The attacker combines data (D) and trigger (T) to build poisoned data (D+T) and release it.
+Victims download the released data and adopt it to train infected DNN models.
+Notably, during the training process, various data transformations/augmentations (F) and hyperparameters (H)†††In this context, H encompasses various training settings, encompassing elements like the model architecture, optimizer choice, and training hyperparameters.
+have been applied by victims.
+Generally, the attacker has no access to any information of the victim phase and only performs the poisoning injection at the attack phase.
+However, proxy attack-based sample selection methods need to select samples with the help of the complete attack process (both attacker and victim phases).
+We argue that all parts involved in backdoor attacks are the factors that affect the efficiency of poisoning samples, containing D, T, F, and H.
+When there contains some gaps (e.g.
+Transform (F) and Hyperparameters (H) in Figure 1) between proxy poisoning attack and actual poisoning process of victims, the poisoning effectiveness of the selected samples may be decreased.
+Following, we will explore what will happen if the actual poisoning process is different from the proxy poisoning attack in proxy attack-based sample selection methods.
+III-A Experimental Settings
+To investigate the proposed question within this section, we engage in a series of experiments utilizing the VGG16 model [43] as the victim model.
+These experiments are conducted on the CIFAR-10 dataset [23].
+Our chosen config encompasses the SGD optimizer, initially set with a learning rate of 0.01, a momentum value of 0.9, and a weight decay coefficient of 5e-4.
+Notably, we incorporate two widely adopted data transformations—random cropping and random horizontal flipping.
 The training is performed for 70 epochs with a batch size of 256.
-Poisoned samples are generated using a blending strategy [4] in which the clean samples are blended with a trigger image using the equation , where is set to 0.15.
-We set the attack target to category 0.
-We adopt the FUS method with a poisoned rate (the size of the poisoning set is ) and the number of selection iterations following [39].
+Poisoning samples are generated using a blending strategy [4] in which the clean samples are blended with a trigger image using the equation , where is set to 0.15.
+We set category 0 as the target for the attack, and we set the poisoning rate with —equating to a poisoning set size of .
+The process involves 15 iterations of selection, aligning with the settings of [50].
 In accordance with previous literature [26], the attack success rate is defined as the probability of classifying a poisoned test data to the target label.
 Unless stated otherwise, all experiments in this section follow the aforementioned settings.
-3.2 Empirical Studies
-Different Data Transforms (F) will decrease the effectiveness of proxy-based sample selection method.
-We investigated how different data transforms (F) affect the effectiveness of proxy-based sample selection methods.
+To explore the degeneration in poisoning efficiency within selected samples due to disparities between the proxy poisoning attack and the actual poisoning process of victims, we compare the Filtering-and-Updating Strategy (FUS) with SELF-FUS.
+SELF-FUS represents an ideal scenario for FUS, assuming the attacker possesses complete information about both the attacker and victim phases during poisoning-based backdoor attacks—a scenario that doesn’t align with the threat model in this study.
+Therefore, proxy poisoning attack of attacker phase is consistent with the actual poisoning process of victim phase in SELF-FUS, which is not guaranteed in the FUS.
+III-B Empirical Studies
+Impact of Different Data Transformations (F) on Proxy Attack-based Sample Selection Efficacy.
+We investigated how different data transforms (F) affect the effectiveness of proxy attack-based sample selection methods.
 To conduct our analysis, we used the experimental settings described previously and followed them exactly.
-Specifically, we set the data transform for the proxy poisoning attack in the sample selection phase (FUS) to ”RandomCrop” and ”RandomHorizontalFlip”, and we used ”None”, ”RandomCrop”, ”RandomHorizontalFlip”, ”RandomRotation”, and ”ColorJitter” for the data transform during the actual poisoning attack in the victim phase.
-Our results, depicted in Figure 2, indicate that different data transforms (F) decrease the effectiveness of proxy-based sample selection methods.
-Different Hyperparameter (H) will decrease the effectiveness of proxy-based sample selection method.
-Furthermore, we explore the impact of different hyperparameter to efficient data searched by FUS [39].
-Following the same experimental settings as described above, we conduct experiments using the hyperparameter ”V-16, SGD, 0.01” to search for efficient sets of backdoor attacks and evaluate the attack success rate on several other hyperparameters, including ”V-13, SGD, 0.01”, ”V-13, Adam, 0.001”, ”V-16, Adam, 0.001”, ”R-18, SGD, 0.01”, ”R-18, Adam, 0.001”, and ”P-18, SGD, 0.01”†††V-13, V-16, R-18, and P-18 denote VGG13, VGG-16, ResNet-18, and PreActResNet-18, respectively.
+Specifically, we configured the data transformations for the proxy poisoning attack during the sample selection phase (FUS) as ”RandomCrop” and ”RandomHorizontalFlip”.
+Subsequently, during the actual poisoning attack within the victim phase, we introduced variations by employing a range of transformations, including ”None”, ”RandomCrop”, ”RandomHorizontalFlip”, ”RandomRotation”, and ”ColorJitter”.
+Our results, depicted in Figure 2, indicate that different data transforms (F) decrease the effectiveness of proxy attack-based sample selection methods.
+An exception to this pattern emerges in cases where the settings for both the proxy and actual attacks align closely, exemplified by the ”1+2” searching and ”1” attacking scenario.
+This particular outcome is logical and can be attributed to the inherent margin of error.
+Effect of Different Hyperparameters (H) on Proxy Attack-based Sample Selection Efficiency.
+In this section, We thoroughly investigated the impact of different hyperparameters on the effective data efficiency searched by FUS [50].
+Adhering to the aforementioned experimental framework, we configured the hyperparameter combination for the proxy poisoning attack during the sample selection phase as ”V-16, SGD, 0.01”.
+Subsequently, during the actual poisoning attack within the victim phase, we introduced variations by employing a range of hyperparameters, including ”V-13, SGD, 0.01”, ”V-13, Adam, 0.001”, ”V-16, Adam, 0.001”, ”R-18, SGD, 0.01”, ”R-18, Adam, 0.001”, and ”P-18, SGD, 0.01”‡‡‡In this context, V-13, V-16, R-18, and P-18 denote VGG13, VGG16, ResNet18, and PreActResNet18, respectively.
 .
-Our results, shown in Figure 3, demonstrate that different hyperparameters (H) can also significantly decrease the effectiveness of the proxy-based sample selection method.
-Overall, our results suggest that the efficacy of proxy-based sample selection methods is sensitive to attacker-agnostic data transforms and hyperparameters used during the victim phase.
-This highlights the importance of using proxy-free search methods to achieve efficient sample selection in backdoor attacks.
+Figure 3 indicates that different hyperparameters (H) decrease the effectiveness of proxy attack-based sample selection methods.
+Once again, we observed anomalous occurrences when the parameters for both proxy searching and actual attacks closely converged, as illustrated by the scenario involving ”V-16-SGD-0.01” proxy searching and subsequent ”V-13-SGD-0.01” attacks within the victim phase.
+Summary Collectively, our findings underscore the susceptibility of proxy attack-based sample selection techniques to shifts in data transforms and hyperparameters deployed during the victim phase.
+Notably, SELF-FUS consistently outperforms FUS across various scenarios, with the exception of outliers observed in both Figure 2 and Figure 3.
+These outliers, marked by the ”1+2” searching and ”1” attacking in Figure 2, as well as the ”V-16-SGD-0.01” searching and ”V-13-SGD-0.01” attacking in Figure 3, share closely aligned proxy and actual attack settings.
+To further validate our hypothesis, we continue to simulate more realistic attack scenarios that introduce disparities in both data transformations (F) and hyperparameters (H) between proxy and actual attacks.
+Figure 4 shows the ASR comparison between FUS and SELF-FUS on both CIFAR-10 and Tiny-ImageNet datasets.
+The notable difference in ASR between FUS and SELF-FUS underscores the significant impact of proxy attack-based methods on sample efficiency.
+Specifically, it highlights how FUS experiences serious degradation in sample efficiency when there is a substantial difference between the settings for proxy poisoning attacks during the attacker phase and those employed in the actual poisoning process during the victim phase.
+IV Methodology
+To improve poisoning efficiency, it is important to develop proxy attack-free sample selection methods that leverage the forensic features of efficient data in poisoning-based backdoor attacks.
+Before introducing our method, we will first discuss some key observations.
+IV-A Forensic Features of Efficient Data in Poisoning-based Backdoor Attacks
+We assert that the primary reason for the efficiency of backdoor injection is due to the similarity between benign and corresponding poisoning samples.
+Typically, poisoning samples exhibit different labels than their clean counterparts, signifying that these pairs share similar attributes yet are categorized differently in dirty-label backdoor attacks.
+In this context, samples exhibiting a high degree of similarity can be considered challenging samples in the context of poisoning tasks.
+Consequently, they tend to be more efficient than low-similarity samples for the purpose of backdoor attacks.
+To substantiate this assumption, we conducted empirical investigations utilizing the CIFAR-10 dataset while adhering to the same experimental settings as detailed in Sec.
+III-A.
+The left segment of Figure 5 visually captures the distribution of similarities among 500 randomly chosen samples and 500 samples searched using the FUS method [50].
+The result demonstrates that the FUS-searched efficient sample exhibits higher similarity compared to randomly selected samples.
+We proceeded to explore the contribution of similarity between benign and corresponding poisoning samples to the efficacy of backdoor attacks.
+The right part of Figure 5 showcases the attack success rate across varying sets of poisoning samples.
+Notably, our observations align with our expectations: using high-similarity samples (Top-500 similarity sampling) for poisoning yielded significantly higher attack success rates than utilizing low-similarity samples (Bottom-500 similarity sampling).
+Furthermore, we compared random sampling with similarity-based selection.
+However, both high-similarity samples (Top-500 similarity sampling) and low-similarity samples (Bottom-500 similarity sampling) for poisoning yielded lower attack success rates compared to random sampling (Random sampling).
+This implies that while similarity can effectively filter out inefficient samples, it lacks a positive correlation with data efficiency in the context of poisoning attacks.
+We argue that the degeneration observed in the performance of the high-similarity samples (Top-500 similarity sampling) can be attributed to the excessively constrained diversity within the pool of poisoning samples.
+The middle part of Figure 5 further illustrates this point by displaying the t-SNE representation of the entire 50,000-sample dataset and the 500 selected samples.
+It is evident that the high-similarity samples exhibit limited diversity.
+To address this limitation, we sampled 500 specimens randomly from the top 5000 samples with the highest similarity (Top-5K similarity random 500 sampling).
+This alleviates the diversity constraint, as confirmed by the right segment of Figure 5.
+Among these approaches, the Top-5K similarity random 500 sampling method is more efficient than random sampling method.
+In summary, the similarity between benign and corresponding poisoning samples and the diversity within the poisoning sample set emerge as two pivotal factors influencing the sample efficiency in backdoor attacks.
+IV-B Threat Model
+Attacker’s Capacities.
+In accordance with the fundamental prerequisites of poisoning-based backdoor attacks [26, 50], we assume that attackers possess the ability to poison a fraction of the training dataset.
+However, they do not possess access to any additional training components encompassed within the victim phase, including the training loss, schedule, or model architecture.
+This assumption describes a more challenging and realistic scenario [26, 46], illustrating the attackers’ constrained knowledge about the target system.
+Furthermore, we also assume that attackers require a pre-trained feature extractor derived from the accessed training dataset.
+This assumption is widely acknowledged within the current backdoor attacks [61, 26].
+Attacker’s Goals.
+Our objective is to augment the effectiveness of dirty-label poisoning backdoor attacks costless, achieved through the judicious selection of a suitable poisoning set from the dataset .
+This approach is orthogonal to most of the current methods for improving poisoning efficiency through designing trigger.
+Therefore, our method can be integrated to other attacks at a fraction of the cost.
 Input:
-Clean training set ; Size of clean training set ; Backdoor trigger ; Attack target ; Poisoned rate ; Diversity rate ; Pre-trained feature extractor ; Trigger function
+Clean training set ; Size of clean training set ; Backdoor trigger ; Attack target ; poisoning rate ; Diversity rate ; Pre-trained feature extractor ; Trigger function
 Output:
 Build the poisoned set ;
-4 Methodology
-To improve poisoning efficiency, it is important to develop proxy-free sample selection methods that leverage the forensic features of efficient data in poisoning-based backdoor attacks.
-Before introducing our method, we will first discuss some key observations.
-4.1 Forensic Features of Efficient Data in Poisoning-based Backdoor Attacks
-We contend that the primary reason for the efficiency of backdoor injection lies in the similarity between clean and corresponding poisoned samples.
-Typically, the labels of poisoned samples are different from those of the original clean samples, meaning that clean and corresponding poisoned samples share similar features but have completely different labels.
-In this case, samples with high similarity can be viewed as challenging samples in the poisoning task, making them more efficient than low-similarity samples for backdoor attacks.
-To validate this hypothesis, we conducted empirical studies on the CIFAR-10 dataset using the same experimental settings as in Sec.
-3.1.
-The left part of Figure 4 depicts the similarity distribution of 500 randomly selected samples and 500 FUS-searched samples [39], which demonstrates that FUS-searched efficient data has higher similarity than random samples.
-We further examined the contribution of similarity to the data efficiency of backdoor attacks by selecting poisoned samples based on similarity.
-The right part of Figure 4 shows the attack success rate when different poisoned samples were used.
-As expected, poisoning with high-similarity samples (Top-500 similarity sampling) resulted in a significantly higher attack success rate than poisoning with low-similarity samples (Bottom-500 similarity sampling).
-Furthermore, we compared random sampling with similarity-selected sampling.
-However, both poisoning with high-similarity samples (Top-500 similarity sampling) and low-similarity samples (Bottom-500 similarity sampling) yielded lower attack success rates than poisoning with random samples (Random sampling).
-This suggests that similarity can only be used to filter out inefficient samples and has no positive correlation with data efficiency in poisoned attacks.
-We argue that the degradation of Top-500 similarity samples can be attributed to the overly limited diversity of the poisoned samples.
-The middle part of Figure 4 shows the t-SNE of the 50K whole dataset and the 500 selected samples, indicating that Top-500 similarity samples have limited diversity.
-Selecting 500 samples randomly from the 5K samples with the highest similarity (Top-5K similarity random 500 sampling) can effectively mitigate this limitation, as shown in the right part of Figure 4.
-Among them, Top-5K similarity random 500 sampling outperforms random sampling and achieves competitive results with only a slight cost compared to FUS [39] methods.
-In summary, the similarity between clean and corresponding poisoned samples and the diversity of poisoned sample set are two key factors in the efficient data of backdoor attacks.
-4.2 Threat Model
-Attacker’s Capacities.
-Following the basic requirement of poisoning-based backdoor attacks [22, 39], we assume that attackers are permitted to poison a portion of the training data, but are not privy to any other training components in the victim phase, such as the training loss, schedule, or model structure.
-These assumptions align with potential real-world scenarios in which such threats can occur.
-Attacker’s Goals.
-Our aim is to enhance the efficiency of poisoning attacks without incurring additional costs, by selecting an appropriate poisoning set from the dataset .
-This approach is distinct from the current methods that enhance efficiency by designing triggers, and hence, our method can be easily incorporated into other attack technologies at minimal cost.
-4.3 Proxy-Free Selection Strategy
-Motivated by the observation of forensic features that contribute to efficient data in poisoning-based backdoor attacks, as demonstrated in Section 4.1, we propose a simple yet effective sample selection strategy (PFS) for enhancing poisoning efficiency.
-Unlike previous studies that rely on a proxy backdoor attack for sample selection, our method is based on the similarity between clean and corresponding poisoned samples, which is only dependent on the Data (D) and Trigger (T), and independent of Data Transform (F) and Hyperparameters (H) (proxy-free).
-Specifically, we utilize a pre-trained feature extractor to compute the cosine similarity (cos()) in the feature space between clean and corresponding poisoned samples.
-We then select the top most similar samples and randomly sample from this set as the poisoning set, where is a hyper-parameter that controls the diversity of the poisoned samples.
-The algorithmic procedure of our method is presented in Algorithm 1.
-Moreover, we demonstrate that our method can filter out most of the worthless samples, which also enhances the FUS search [39].
-Therefore, we propose a combined approach (FUS+PFS) that utilizes the FUS search on the coarse poisoned set containing samples to further enhance the poisoning efficiency.
-5 Experiments
+IV-C Proxy Attack-free Selection Strategy
+Drawing inspiration from our analysis of forensic features contributing to efficient data in poisoning-based backdoor attacks, as showcased in Sec.
+IV-A, we introduce a simple yet efficient sample selection strategy termed the Proxy attack-Free Strategy (PFS) for improving poisoning efficiency.
+Unlike prior approaches that rely on a proxy attack task for sample selection, our technique capitalizes on the similarity between benign and corresponding poisoning samples, which is only related to the Data (D) and Trigger (T), liberating it from dependence on the settings of actual attack within victim phase.
+Specifically, we adopt a pre-trained feature extractor to compute the cosine similarity (cos()) within the feature space between benign and corresponding poisoning samples.
+Subsequently, we identify the top most similar samples and randomly select from this subset as the poisoning set, where is a hyper-parameter that modulates the diversity of the poisoning samples.
+The procedure of our methodology is presented in Algorithm 1.
+Acknowledging that our PFS relies on a feature extractor pre-trained on a clean dataset to compute similarity, we contend that the essence of FUS and PFS diverges significantly.
+Our PFS identify the forensic features of efficient samples in poisoning-based backdoor attacks, and adopt the similarity and diversity to identify efficient samples, which is disentangled from the actual attack within victim phase.
+Therefore, the effectiveness of PFS primarily depends on the accuracy of similarity measurement.
+Furthermore, we confirm that our method efficiently filters out most of the ineffective samples, thus enhancing the performance of FUS [50].
+Therefore, we introduce a collaborative strategy called FUS+PFS, which combines FUS search with the initial coarse poison set containing samples to further enhance poisoning efficiency.
+IV-D Theoretical Analyses
+Additionally, we provide theoretical analyses of the proposed PFS in terms of both active learning and neural tangent kernel.
+Perspective of Active Learning.
+The process of selecting efficient poisoning samples in the context of backdoor attacks can be analogized to the principles of active learning.
+Active learning (AL) is a machine learning paradigm that efficiently acquires annotated data from an extensive pool of unlabeled data [35].
+Its objective is to prioritize human labeling efforts on the most informative data points, leading to improved model performance while reducing data annotation costs.
+Similarly, the task of selecting efficient poisoning samples in a backdoor attack aims to identify the most informative data points that can optimize the efficiency of the poisoning process, thereby minimizing the costs for the attacker.
+A pivotal aspect of active learning encompasses the concepts of uncertainty and diversity in the dataset [6].
+Methods that rely on uncertainty utilize the predictive confidence of model to select challenging examples, while diversity-based sampling leverages the dissimilarity among data points, often involving clustering techniques.
+Notably, the combination of uncertainty and diversity strategies has demonstrated significant success in active learning [56, 42].
+Similarly, our proposed PFS takes into account both uncertainty (Line 8 in Algorithm 1) and diversity (Line 9 in Algorithm 1) of data points.
+Furthering our understanding of data point uncertainty within backdoor attacks, we delve into more comprehensive analyses to underscore the connection between the proposed measure of similarity between benign and corresponding poisoning samples and the expression of data point uncertainty.
+Drawing from theoretical insights provided by margin theory in active learning [2], we recognize that examples near the decision boundary offer significant potential for reducing annotation requirements.
+In the context of a backdoor attack, we exclusively consider the true class and the attack-target class of the clean sample and its corresponding poisoning sample , respectively, forming a two-class classification task.
+For any given poisoning sample , its corresponding clean sample is the nearest sample, albeit labeled with a different category.
+This configuration establishes a margin , with the decision boundary residing centrally within the margin.
+As such, the distance between the sample and the decision boundary is proportionate to .
+To summarize, the distance between benign and corresponding poisoning samples, , can be employed to convey the uncertainty of data points.
+This assessment is carried out using a pre-trained feature extractor to compute distances within the feature space.
+Figure 6 illustrates the effective integration of individual similarity and set diversity in constructing a poisoning injection.
+Perspective of Neural Tangent Kernel (NTK).
+To further explain this interesting phenomenon (i.e., similarity between benign and corresponding poisoning samples and the diversity within the poisoning sample set emerge as two pivotal factors influencing the efficiency of data in backdoor attacks), we exploit recent studies on NTK [22, 15, 16, 30] for analyzing the sample efficiency of backdoor attacks:
+Theorem 1.
+Suppose the training dataset consists of benign samples and poisoned samples , whose images are i.i.d.
+sampled from uniform distribution and belonging to classes.
+Assume that the DNN is a multivariate kernel regression and is trained via .
+For the expected predictive confidences over the target label , we have: , where is poisoning testing samples of attacks.
+In general, Theorem 1 suggests that selecting samples characterized by high similarity between clean () and corresponding poisoning samples (), as well as high similarity between poisoning samples () and poisoned testing samples (), enhances the confidence in predicting poisoned samples to the target class.
+In the context of poisoning-based backdoor attacks, where training samples share the same distribution as testing samples, high similarity between poisoning samples and poisoned testing samples indicates congruence in the distribution between the poisoning set and the training set, signifying a high level of diversity within the poisoning set.
+The proof of this theorem can be found in the Sec.
+A of the Supplementary Materials.
+V Experiments
 We evaluate the effectiveness of PFS on datasets: CIFAR -10 (Sec.
-5.2), CIFAR-100 (Sec.
-5.3), and Tiny-ImageNet (Sec.
-A.4 of the Supplementary Materials).
+V-B), Tiny-ImageNet (Sec.
+V-C), and CIFAR-100 (Sec.
+E of the Supplementary Materials).
 We also provide experimental settings in Sec.
-5.1, time-consuming analysis in Sec.
-5.4, and ablation study in Sec.
-5.5.
-5.1 Experimental Settings
-We conduct experiments with three trigger types (BadNets [13], Blended [4], Optimized [48]), six data transforms (None, RandomCrop, RandomHorizontalFlip, RandomRotation, ColorJitter, RandomCrop+RandomHorizontalFlip), and three DNN architectures (VGG16 [34], ResNet18 [15], PreAct-ResNet18 [16]).
-We also use four optimizers: SGD-0.01‡‡‡SGD-0.01 means adopting the SGD optimizer with an initial learning rate of 0.01,, SGD-0.02, Adam-0.001, Adam-0.002.
-The adopted feature extractor is the ResNet-18 that has been pre-trained on the current dataset.
-Additionally, in Section A.5 of the Supplementary Materials, we present experiments with different pre-trained feature extractors on the CIFAR-10 dataset.
-Our results show that even using a common feature extractor pre-trained on the ImageNet dataset yields significant improvements compared to the baseline.
-We compare our method to the commonly used random selection (Random) [13, 4] and FUS [39].
-In the FUS search process, we use VGG16, RandomCrop+RandomHorizontalFlip, and SGD-0.01 to construct the proxy attack.
-For our method, we use pre-trained ResNet models for similarity measurement and set the diversity hyperparameter to 10 for all three datasets.
-Other hyperparameters are kept the same as in Sec.
-3.1.
+V-A, comparison with recent sample selection methods in Sec.
+V-D, time consumption analysis in Sec.
+V-F, ablation study in Sec.
+V-G, and experiment against defense method in Sec.
+V-H.
 | Trigger | Data Transform | Model | |||||||||||
@@
 | Avg | 0.903 | 0.969 | 0.946 | 0.975 | 0.893 | 0.966 | 0.938 | 0.975 | 0.927 | 0.986 | 0.966 | 0.990 | |
-5.2 Experiments on CIFAR-10 Dataset
+V-A Experimental Settings
+We conduct comprehensive experiments encompassing three trigger types§§§Additionally, we consider the recent backdoor attacks such as WaNet [38] and ISSBA [27] in Sec.
+V-D.
+(BadNets [14], Blended [4], Optimized [61]), six data transformations (None, RandomCrop, RandomHorizontalFlip, RandomRotation, ColorJitter, RandomCrop+RandomHorizontalFlip), and three different DNN architectures (VGG16 [43], ResNet18 [19], PreAct-ResNet18 [20]).
+Furthermore, we explore the effects of four optimizers: SGD-0.01¶¶¶SGD-0.01 implies the adoption of the SGD optimizer with an initial learning rate of 0.01., SGD-0.02, Adam-0.001, and Adam-0.002.
+The employed feature extractor is ResNet18, pre-trained on the specific dataset using the SGD optimizer with an initial learning rate of 0.01, momentum of 0.9, and weight decay of 5e-4, along with two standard data augmentations: random crop and random horizontal flip.
+Each training iteration spans 70 epochs, conducted with a batch size of 256.
+In addition, Sec.
+V-E presents experiments involving various pre-trained feature extractors on the CIFAR-10 dataset.
+Our findings reveal that even utilizing a commonly used ImageNet pre-trained feature extractor yields notable enhancements compared to the baseline.
+We perform a comprehensive comparative analysis, evaluating our method against the widely employed random selection (Random) [14, 4] and FUS [50] approaches.
+In the FUS search procedure, we utilize the VGG16 architecture, coupled with RandomCrop+RandomHorizontalFlip data transformations and SGD-0.01 as the optimizer to establish the proxy attack.
+Conversely, for our proposed method, we utilize pre-trained ResNet models for similarity measurement and set the diversity hyperparameter to 10 across all three datasets.
+All other hyperparameters remain consistent with those outlined in Sec.
+III-A.
+Additionally, Sec.
+V-D shows the comparison with recent sample selection methods (FUS [50], OPS [17], Gradient[10], RD Score[49]) on the Tiny-ImageNet dataset.
+V-B Experiments on CIFAR-10 Dataset
 Results on different data transforms and architectures.
-Table 1 presents the attack success rate for BadNets, Blended, and Optimized triggers using different data transforms and architectures.
-We use VGG16, RandomCrop+RandomHorizontalFlip, and SGD-0.01 to construct the proxy attack.
-Our PFS method significantly improves the efficiency of poisoned data for backdoor attacks.
-In all 54 settings, the attack success rate of the poisoned samples selected by our method is higher than that of randomly selected samples, at the same poisoning ratios.
-On average, our method outperforms random selection by 0.044, 0.074, and 0.066 for BadNets, Blended, and Optimized triggers, respectively.
-In general, our method performs better than FUS, achieving stronger attack strength in 48 out of 54 settings.
-It is worth noting that the FUS search requires hundreds or even thousands of times more time than our method.
-Combining FUS with our method yields the best results in 51 out of 54 settings, as our method filters out many inefficient samples and narrows the FUS search space.
-Results on different optimizer and architectures.
-Table 2 shows the attack success rate for the BadNets, Blended, and Optimized triggers with different optimizers and architectures.
-We use VGG16, RandomCrop+RandomHorizontalFlip, and SGD-0.01 to construct the proxy attack.
-The results demonstrate the effectiveness of our PFS method.
-First, for all 36 settings, the attack success rate achieved by our method is higher than that achieved by random selection, with average improvements of 0.036, 0.0437, and 0.017 for the BadNets, Blended, and Optimized triggers, respectively.
-Second, our method outperforms FUS in 25 out of 36 cases.
-When combining our method with FUS, we obtain the best results in 33 out of 36 settings.
-In the Supplementary Materials, we have included additional experiments to further evaluate the effectiveness and robustness of our proposed strategy.
-Specifically, in Sec.
-A.1 and A.2, we present the attack success rate (ASR) and benign accuracy (BA) under poisoned rates of and on the CIFAR-10 dataset.
-The results demonstrate that our method remains robust to different poisoned rates and target labels while preserving benign accuracy.
-Additionally, in Sec.
-A.3, we present the ASR for an additional target class (class 3) on the CIFAR-10 dataset.
-Overall, our findings confirm the effectiveness of our proposed method and support our hypothesis that the efficiency of backdoor attacks is determined by sample similarity and diversity.
+Table I provides a comprehensive overview of the attack success rates for BadNets, Blended, and Optimized triggers, considering various data transforms and architectures.
+Constructing the proxy attack for FUS, we employ VGG16 with RandomCrop+RandomHorizontalFlip and SGD-0.01 as the optimizer.
+Our proposed PFS method significantly enhances the poisoning efficacy in backdoor attacks.
+Across all 54 configurations, the attack success rate of poisoning samples selected through our approach consistently surpasses that of randomly selected samples at identical poisoning rates.
+On average, our method achieves a improvement of 0.044, 0.074, and 0.066 in terms of attack success rate for BadNets, Blended, and Optimized triggers, respectively.
+In general, our method outperforms FUS, demonstrating greater ASR in 48 of the 54 scenarios.
+Notably, FUS searches necessitate significantly more time compared to our method, often by hundreds or even thousands of times.
+Furthermore, when the model is VGG16 and the trigger is set to Blended, FUS surpasses our PFS.
+This can be attributed to two factors: (i) Both the proxy attack model employed in FUS and the model utilized in the victim phase are VGG16; (ii) In comparison to BadNets and Optimized triggers, the Blended trigger increases the similarity distance between benign and corresponding poisoning samples, thereby marginally impacting the performance of our proposed PFS.
+The combination of FUS and PFS, denoted as FUS+PFS, excels in 51 out of the 54 scenarios.
+This outcome stems from PFS effectively eliminating inefficient samples, consequently decreasing the search space for FUS.
 | Trigger | Optimizer | Model | |||||||||||
@@
 | Adam-0.001 | 0.886 | 0.910 | 0.937 | 0.909 | 0.900 | 0.953 | 0.951 | 0.970 | 0.896 | 0.949 | 0.943 | 0.962 | |
-| Avg | 0.881 | 0.891 | 0.927 | 0.921 | 0.883 | 0.944 | 0.947 | 0.968 | 0.884 | 0.944 | 0.939 | 0.963 | |
+| Avg | 0.856 | 0.891 | 0.927 | 0.921 | 0.883 | 0.944 | 0.947 | 0.968 | 0.884 | 0.944 | 0.939 | 0.963 | |
 | Optimized | SGD-0.01 | 0.979 | 0.999 | 0.996 | 1.000 | 0.979 | 0.999 | 0.997 | 1.000 | 0.982 | 0.999 | 0.997 | 1.000 |
@@
 | Avg | 0.978 | 0.999 | 0.997 | 1.000 | 0.983 | 0.999 | 0.997 | 1.000 | 0.983 | 0.997 | 0.996 | 1.000 | |
+Results on different optimizer and architectures.
+Table II presents the ASR for the BadNets, Blended, and Optimized triggers, considering different optimizers and architectures.
+We utilize VGG16 with RandomCrop+RandomHorizontalFlip data transformations and SGD-0.01 as the optimizer for constructing the proxy attack in FUS.
+The outcomes underscore the effectiveness of our PFS method.
+Firstly, across all 36 configurations, our PFS consistently achieves higher attack success rates than random selection, with average improvements of 0.0437, 0.052 and 0.017 for the BadNets, Blended, and Optimized triggers, respectively.
+Secondly, our PFS outperforms FUS in 25 out of the 36 cases.
+Similar to previous experiments, when the model is VGG16 and the trigger is set to Blended, FUS outperforms our PFS.
+Notably, the combination of our method with FUS, referred to as FUS+PFS, achieves the best performance in 33 out of the 36 settings.
+The Supplementary Materials contain additional experiments aimed at further evaluating the effectiveness and robustness of our proposed strategy.
+Specifically, in Sec.
+B and Sec.
+C, we provide the attack success rate (ASR) and benign accuracy (BA) under poisoning rates of and on the CIFAR-10 dataset, respectively.
+The outcomes illustrate the robustness of our method across different poisoning rates while maintaining benign accuracy.
+Moreover, in Sec.
+D, we present the ASR for an additional target class (class 3) on the CIFAR-10 dataset.
+Overall, our results consistently affirm the effectiveness of the proposed pfs and support our hypothesis that the efficiency of backdoor attacks depends on the interaction between individual similarity and set diversity.
+V-C Experiments on Tiny-ImageNet Dataset
+Results on different data transforms and architectures.
+Table III shows that PFS outperforms the random method in terms of attack success rates of poisoning samples on Tiny-ImageNet, with average boosts of 0.054, 0.094, and 0.072 on BadNets, Blended, and Optimized triggers, respectively.
+In comparison to FUS, PFS outperforms FUS in 45 out of 54 cases.
+In contrast to experiments on the CIFAR-10 dataset, our proposed PFS method consistently achieves higher average ASR than FUS across different data transformations on the Tiny-ImageNet dataset.
+When the two methods are combined, the best results are achieved in 34 out of 54 settings, indicating the superiority of our method.
+These findings provide strong evidence supporting the effectiveness of our proposed method on the scenarios involving extensive datasets.
+Results on different data transforms, optimizer, and architectures.
+To further illustrate the effectiveness of our proposed PFS, Figure 7 visually shows the average ASR across various data transforms and optimizers on the Tiny-ImageNet dataset.
+The findings indicate that all advanced sample selection strategies (FUS, PFS, and FUS+PFS) achieve notably superior results compared to the random selection strategy (RSS).
+Furthermore, our proposed PFS consistently achieves better ASR compared to FUS across all settings.
 | Trigger | Data Transform | Model | |||||||||||
@@
 | Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | ||
-| BadNets | None | 0.933 | 0.970 | 0.965 | 0.971 | 0.950 | 0.978 | 0.973 | 0.979 | 0.949 | 0.977 | 0.970 | 0.977 |
-| RandomCrop | 0.882 | 0.947 | 0.930 | 0.945 | 0.894 | 0.951 | 0.940 | 0.952 | 0.888 | 0.951 | 0.940 | 0.951 | |
-| RandomHorizontalFlip | 0.920 | 0.964 | 0.958 | 0.965 | 0.936 | 0.970 | 0.965 | 0.971 | 0.929 | 0.968 | 0.962 | 0.968 | |
-| RandomRotation | 0.868 | 0.923 | 0.912 | 0.924 | 0.866 | 0.939 | 0.925 | 0.938 | 0.875 | 0.935 | 0.926 | 0.938 | |
-| ColorJitter | 0.934 | 0.970 | 0.965 | 0.971 | 0.952 | 0.978 | 0.973 | 0.979 | 0.946 | 0.977 | 0.970 | 0.978 | |
-| RandomCrop+ RandomHorizontalFlip | 0.879 | 0.939 | 0.929 | 0.939 | 0.888 | 0.947 | 0.942 | 0.950 | 0.894 | 0.949 | 0.941 | 0.951 | |
-| Avg | 0.903 | 0.952 | 0.943 | 0.953 | 0.914 | 0.961 | 0.953 | 0.962 | 0.914 | 0.960 | 0.952 | 0.961 | |
-| Blended | None | 0.866 | 0.923 | 0.920 | 0.941 | 0.884 | 0.963 | 0.946 | 0.974 | 0.894 | 0.975 | 0.950 | 0.982 |
-| RandomCrop | 0.863 | 0.874 | 0.891 | 0.880 | 0.826 | 0.916 | 0.906 | 0.938 | 0.829 | 0.931 | 0.912 | 0.942 | |
-| RandomHorizontalFlip | 0.891 | 0.945 | 0.935 | 0.960 | 0.902 | 0.968 | 0.951 | 0.977 | 0.912 | 0.974 | 0.958 | 0.982 | |
-| RandomRotation | 0.818 | 0.862 | 0.901 | 0.865 | 0.841 | 0.916 | 0.916 | 0.936 | 0.848 | 0.926 | 0.921 | 0.944 | |
-| ColorJitter | 0.863 | 0.920 | 0.917 | 0.931 | 0.904 | 0.972 | 0.956 | 0.977 | 0.913 | 0.979 | 0.959 | 0.983 | |
-| RandomCrop+ RandomHorizontalFlip | 0.811 | 0.865 | 0.905 | 0.872 | 0.826 | 0.918 | 0.920 | 0.942 | 0.850 | 0.931 | 0.924 | 0.947 | |
-| Avg | 0.852 | 0.898 | 0.912 | 0.908 | 0.864 | 0.942 | 0.933 | 0.957 | 0.874 | 0.953 | 0.937 | 0.963 | |
-| Optimized | None | 0.979 | 0.999 | 0.995 | 0.999 | 0.884 | 0.972 | 0.931 | 0.982 | 0.895 | 0.983 | 0.962 | 0.987 |
-| RandomCrop | 0.989 | 0.999 | 0.997 | 1.000 | 0.983 | 0.999 | 0.996 | 1.000 | 0.981 | 0.999 | 0.995 | 1.000 | |
-| RandomHorizontalFlip | 0.985 | 0.999 | 0.997 | 1.000 | 0.938 | 0.998 | 0.980 | 0.983 | 0.938 | 0.997 | 0.984 | 0.998 | |
-| RandomRotation | 0.810 | 0.831 | 0.867 | 0.785 | 0.827 | 0.909 | 0.882 | 0.882 | 0.835 | 0.908 | 0.883 | 0.900 | |
-| ColorJitter | 0.984 | 0.997 | 0.993 | 1.000 | 0.967 | 0.989 | 0.979 | 0.985 | 0.956 | 0.993 | 0.990 | 0.994 | |
-| RandomCrop+ RandomHorizontalFlip | 0.991 | 0.999 | 0.997 | 1.000 | 0.987 | 0.999 | 0.998 | 1.000 | 0.982 | 0.998 | 0.994 | 0.999 | |
-| Avg | 0.956 | 0.971 | 0.974 | 0.964 | 0.931 | 0.978 | 0.961 | 0.972 | 0.931 | 0.980 | 0.968 | 0.980 | |
-| PFS (w/o pre-training) | PFS (w pre-training) | FUS () |
-| 17 | 562 | 9,150 |
-5.3 Experiments on CIFAR-100 Dataset
-The results for CIFAR-100 in Table 3 are similar to those for CIFAR-10.
-Our proxy-free approach outperforms random selection in all cases, with average boosts of 0.047, 0.068, and 0.037 for BadNets, Blended, and Optimized triggers, respectively.
-Compared to FUS, PFS achieves better attack results in 49 out of 54 settings.
-5.4 Analysis of Time Consumption
-Table 4 presents a comparison of the running times of our PFS with those of FUS.
-Our method demonstrates remarkable speed, being faster than FUS.
-Even with pre-training time included, our method is still approximately faster.
+| BadNets | None | 0.952 | 0.990 | 0.983 | 0.989 | 0.960 | 0.992 | 0.985 | 0.991 | 0.956 | 0.991 | 0.981 | 0.989 |
+| RandomCrop | 0.893 | 0.959 | 0.950 | 0.958 | 0.917 | 0.970 | 0.963 | 0.973 | 0.904 | 0.973 | 0.958 | 0.971 | |
+| RandomHorizontalFlip | 0.952 | 0.989 | 0.984 | 0.989 | 0.963 | 0.991 | 0.981 | 0.989 | 0.947 | 0.983 | 0.974 | 0.985 | |
+| RandomRotation | 0.881 | 0.932 | 0.937 | 0.944 | 0.912 | 0.959 | 0.964 | 0.966 | 0.910 | 0.957 | 0.963 | 0.970 | |
+| ColorJitter | 0.956 | 0.990 | 0.985 | 0.990 | 0.966 | 0.993 | 0.986 | 0.992 | 0.956 | 0.990 | 0.983 | 0.988 | |
+| RandomCrop+ RandomHorizontalFlip | 0.900 | 0.943 | 0.955 | 0.950 | 0.917 | 0.977 | 0.967 | 0.975 | 0.924 | 0.974 | 0.965 | 0.978 | |
+| Avg | 0.922 | 0.967 | 0.966 | 0.970 | 0.939 | 0.980 | 0.974 | 0.981 | 0.933 | 0.978 | 0.971 | 0.980 | |
+| Blended | None | 0.889 | 0.964 | 0.947 | 0.975 | 0.876 | 0.965 | 0.930 | 0.973 | 0.873 | 0.969 | 0.941 | 0.978 |
+| RandomCrop | 0.836 | 0.928 | 0.920 | 0.942 | 0.835 | 0.931 | 0.914 | 0.936 | 0.822 | 0.927 | 0.906 | 0.932 | |
+| RandomHorizontalFlip | 0.905 | 0.966 | 0.964 | 0.980 | 0.888 | 0.964 | 0.950 | 0.977 | 0.886 | 0.965 | 0.954 | 0.981 | |
+| RandomRotation | 0.865 | 0.931 | 0.946 | 0.953 | 0.872 | 0.945 | 0.942 | 0.966 | 0.870 | 0.950 | 0.941 | 0.966 | |
+| ColorJitter | 0.883 | 0.958 | 0.945 | 0.964 | 0.871 | 0.956 | 0.917 | 0.962 | 0.872 | 0.967 | 0.931 | 0.971 | |
+| RandomCrop+ RandomHorizontalFlip | 0.850 | 0.931 | 0.940 | 0.958 | 0.840 | 0.929 | 0.919 | 0.942 | 0.828 | 0.931 | 0.912 | 0.940 | |
+| Avg | 0.871 | 0.946 | 0.944 | 0.962 | 0.864 | 0.948 | 0.929 | 0.959 | 0.859 | 0.952 | 0.931 | 0.961 | |
+| Optimized | None | 0.907 | 0.981 | 0.956 | 0.984 | 0.875 | 0.968 | 0.934 | 0.968 | 0.862 | 0.970 | 0.935 | 0.967 |
+| RandomCrop | 0.897 | 0.976 | 0.959 | 0.980 | 0.886 | 0.968 | 0.950 | 0.972 | 0.869 | 0.958 | 0.945 | 0.951 | |
+| RandomHorizontalFlip | 0.903 | 0.979 | 0.956 | 0.985 | 0.892 | 0.976 | 0.950 | 0.976 | 0.877 | 0.975 | 0.948 | 0.974 | |
+| RandomRotation | 0.899 | 0.977 | 0.957 | 0.972 | 0.897 | 0.978 | 0.956 | 0.972 | 0.888 | 0.973 | 0.948 | 0.956 | |
+| ColorJitter | 0.695 | 0.714 | 0.744 | 0.707 | 0.769 | 0.773 | 0.803 | 0.775 | 0.772 | 0.779 | 0.816 | 0.781 | |
+| RandomCrop+ RandomHorizontalFlip | 0.892 | 0.973 | 0.960 | 0.974 | 0.890 | 0.964 | 0.952 | 0.968 | 0.870 | 0.963 | 0.949 | 0.957 | |
+| Avg | 0.866 | 0.933 | 0.922 | 0.934 | 0.868 | 0.938 | 0.924 | 0.939 | 0.856 | 0.936 | 0.924 | 0.931 | |
+| Dataset | PFS (w/o pre-training) | PFS (w pre-training) | FUS |
+| CIFAR-10 | 17 | 562 | 9,150 |
+| Tiny-ImageNet | 19 | 6784 | 106284 |
+V-D Comparison with Recent Sample Selection Strategies
+Table IV presents a comparative analysis between our PFS and recent sample selection strategies (FUS [50], OPS [17], Gradient[10], and RD Score[49]) on the Tiny-ImageNet dataset.
+Various recent triggers (Blended, WaNet, and ISSBA) are considered in this experiment.
+The results demonstrate that all sample selection strategies exhibit superior ASR compared to the random selection strategy (RSS), with our proposed PFS consistently demonstrating the highest sample efficiency across all settings when compared to other recent sample selection strategies.
+V-E Experiments on Different Pre-trained Feature Extractor
+Although both FUS and PFS rely on a proxy task, there are essential differences between proxy tasks.
+FUS relies on a proxy attack task closely resembling the actual attack.
+Nevertheless, due to the inherent shortcuts of backdoor learning, overfitting to the data transformation and hyperparameter of proxy attack significantly impact actual injection efficiency.
+In contrast, our PFS selects efficient samples considering individual similarity and set diversity.
+As a result, PFS operates independently of the specific attack deployed during the victim phase, with minimal impact from data transformation and hyperparameters within the pre-trained feature extractor on sample efficiency during this phase.
+This independence is evidenced in our experiments with various pre-trained extractors, as illustrated in Figure 8.
+Remarkably, even a common extractor pre-trained on ImageNet results in considerable improvements compared to the baseline.
+V-F Analysis of Time Consumption
+Table V presents a comparison of the running times of our PFS with those of FUS.
+PFS demonstrates remarkable speed, being faster than FUS on CIFAR-10 dataset.
+Even with pre-training time included, our method is still approximately faster on CIFAR-10 dataset.
 Importantly, our method utilizes the same pre-trained model for different triggers, while FUS needs to conduct a complete search from the beginning for each trigger.
 This advantage makes our method a practical choice for situations where efficiency is crucial.
-5.5 Ablation Study
-We conduct an ablation study on the hyperparameter , which determines the diversity of the poisoned sample set.
+V-G Ablation Study
+We conduct an ablation study on the hyperparameter , which determines the diversity of the poisoning sample set.
 The results are shown in Figure 9, and the attack success rates demonstrate similar trends under different data transforms.
@@
 However, too large of a value for can reduce the role of similarity and weaken the attack.
-These findings highlight the importance of balancing similarity and diversity for efficient poisoned samples, supporting our claim in Sec.
-4.1.
-6 Conclusion and Future Work
+These findings highlight the importance of balancing similarity and diversity for efficient poisoning samples, supporting our claim in Sec.
+IV-A.
+In our PFS, we set in all experimental settings.
+V-H Experiment with Backdoor Defense.
+We evaluate the robustness of PFS against backdoor defense using the pruning-based defense method [32] and the tune-based defense method FST [36].
+Pruning [32] is a widely used defense mechanism.
+It achieves this by pruning neurons that remain inactive in response to benign inputs, thereby neutralizing the effectiveness of the backdoor behavior.
+On the other hand, FST [36] is a tuning-based backdoor purification method.
+It operates by inducing feature shifts through deliberate adjustments to the classifier weights, effectively diverging them from the initially compromised weights.
+As depicted in Figure 10, PFS demonstrates superior robustness against Pruning [32] compared to RSS and FUS.
+Similarly, Figure 11 illustrates that our proposed PFS effectively against the tune-based defense method FST [36] compared to RSS and FUS.
+It is noteworthy that while FST [36] demonstrates superior defense capabilities compared to Pruning [32], it comes at the expense of a substantial reduction in benign accuracy.
+VI Conclusion and Future Work
 Conclusion.
-This paper presents empirical evidence and reveals two intriguing phenomena related to data efficiency in backdoor attacks.
-First, we demonstrate that sample selection methods based on proxy attacks suffer from performance degradation when gaps exist between proxy poisoning attacks and actual victim poisoning processes.
-Second, we show that the similarity between clean and corresponding poisoned samples, as well as the diversity of the poisoned sample set, are key features for efficient data in backdoor attacks.
-Based on these findings, we propose a simple yet effective proxy-free sample selection strategy that selects poisoned samples based on the similarity between clean and corresponding poisoned samples.
-Our method significantly improves the poisoning efficiency on benchmark datasets with almost no additional cost.
+This paper contributes empirical insights into two interesting phenomena associated with data efficiency in backdoor attacks.
+Firstly, we highlight the performance degeneration experienced by sample selection methods relying on proxy attacks when discrepancies arise between proxy attacks within attacker phase and actual poisoning processes within the victim phase.
+Furthermore, we argue that the significance of both the similarity between benign and corresponding poisoning samples and the diversity within the poisoning sample set as essential factors for efficient data in backdoor attacks.
+Leveraging these observations, we introduce a simple yet efficient proxy attack-free sample selection strategy that is driven by the similarity between benign and corresponding poisoning samples.
+Our method significantly improves poisoning efficiency across benchmark datasets with minimal additional cost.
 Future Work.
-While the proxy-free sampling strategy we proposed has demonstrated efficient data selection in backdoor attacks, we note that the similarity between clean and corresponding poisoned samples does not necessarily guarantee high efficiency, as similarity only filters out non-efficient samples.
-In the future, it would be desirable to identify a unified indicator that positively correlates with data efficiency in backdoor attacks.
+While the proposed PFS has demonstrated efficient data selection in backdoor attacks, we note that similarity between benign and corresponding poisoned samples does not necessarily guarantee efficiency, as similarity only filters out non-effective samples.
+In the future, it would be desirable to identify an unified indicator that positively correlates with data efficiency in backdoor attacks.
+Moreover, our proposed PFS is customized specifically for dirty-label backdoor attacks.
+In future research, we intend to explore efficient sample selection strategies designed specifically for clean-label backdoor attacks.
+Acknowledgments
+The work was supported in part by the National Natural Science Foundation of China under Grands U19B2044 and 61836011.
 Appendix A Supplementary Materials
 This supplementary material is organized as follows.
-We first discuss the attack success rate and benign accuracy under poisoned rates with and on the CIFAR-10 dataset in Sec.
-A.1 and Sec.
-A.2, respectively.
-Additionally, we presents the results of attack success rate under additional target classes on the CIFAR-10 dataset in Sec.
-A.3.
+We first discuss the attack success rate and benign accuracy under poisoning rates with and on the CIFAR-10 dataset in Sec.
+A-B and Sec.
+A-C, respectively.
+Additionally, we present the results of attack success rate under additional target classes on the CIFAR-10 dataset in Sec.
+A-D.
 Besides, Sec.
-A.4 presents the experimental results on the Tiny-ImageNet dataset.
-At last, Sec.
-A.5 shows the experiments on different pre-trained feature extractor .
-A.1 Attack Success Rate Under Additional Poisoned Rates on the CIFAR-10 Dataset
-We evaluate the attack success rate (ASR) under different poisoned rates on the CIFAR-10 dataset.
-Tables 5 and 6 present the ASR with trigger BadNets and Blended at poisoned rates of and , respectively.
-Results under poisoned rates of and are similar to those at .
-(i) Experiments under poisoned rates of : Our proxy-free approach outperforms random selection in all cases, with average boosts of 0.030 and 0.055 on BadNets and Blended triggers, respectively.
-The proxy-free method also achieves better attack results than FUS in most cases (27/36).
+V-C presents the experimental results on the Tiny-ImageNet dataset.
+A-A The Proof of Theorem 1
+Theorem 2.
+Suppose the training dataset consists of benign samples and poisoned samples , whose images are i.i.d.
+sampled from uniform distribution and belonging to classes.
+Assume that the DNN is a multivariate kernel regression and is trained via .
+For the expected predictive confidences over the target label , we have: , where is poisoned testing samples of attacks.
+Proof 1 (Proof of Theorem 1:).
+We treat our model as a m-way kernel least square classifier and use a cross-entropy loss for training the kernel.
+The output of is a m-dimensional vector.
+Let us assume be expected predictive confidences corresponding to the target class .
+Following previous works [15, 16], we know the kernel regression solution for neural tangent kernel (NTK) [22] is:
+where is the RBF kernel and .
+Since the training samples are evenly distributed, there are benign samples belonging to .
+Without loss of generality, we assume the target label while others are .
+Then the regression solution can be re-formulated to:
+Therefore, following [30], we have:
+Similar to [16], we also remove the term because typically don’t belong to the target and , otherwise the attacker has no incentive to craft poisoned samples.
+When close to , which implies that the poisoning rate close to , the attacker can achieve the optimal attack efficacy [14, 27]. Given , we have:
+we here let in , because typically be similar as and . Therefore, we have:
+A-B Attack Success Rate Under Additional poisoning rates on the CIFAR-10 Dataset
+We evaluate the attack success rate (ASR) under different poisoning rates on the CIFAR-10 dataset.
+Tables VI and VII present the ASR with trigger BadNets and Blended at poisoning rates of and , respectively.
+Results under poisoning rates of and are similar to those at .
+(i) Experiments under poisoning rates of : Our PFS approach outperforms random selection in all cases, with average boosts of 0.030 and 0.055 on BadNets and Blended triggers, respectively.
+The PFS method also achieves better attack results than FUS in most cases (27/36).
 The combination of these two methods achieves superior results in all 36 settings.
-(ii) Experiments under poisoned rates of : Our proxy-free approach outperforms random selection in all cases, with average boosts of 0.023 and 0.04 on BadNets and Blended triggers, respectively.
-The proxy-free method also achieves better attack results than FUS in most cases (26/36).
+(ii) Experiments under poisoning rates of : Our PFS approach outperforms random selection in all cases, with average boosts of 0.023 and 0.04 on BadNets and Blended triggers, respectively.
+The PFS method also achieves better attack results than FUS in most cases (26/36).
 The combination of these two methods achieves superior results in all 36 settings.
@@
 | Avg | 0.910 | 0.944 | 0.960 | 0.969 | 0.920 | 0.966 | 0.962 | 0.969 | 0.932 | 0.972 | 0.969 | 0.987 | |
-A.2 Benign Accuracy Under Additional Poisoned Rates on the CIFAR-10 Dataset
-We also present the benign accuracy (BA) (i.e., the probability of classifying a benign test data to the correct label) under other poisoned rates on the CIFAR-10 dataset.
-Table 7 and Table 8 illustrate the benign accuracy with trigger BadNets and Blended under poisoned rates of and , respectively.
+A-C Benign Accuracy Under Additional poisoning rates on the CIFAR-10 Dataset
+We also present the benign accuracy (BA) (i.e., the probability of classifying a benign test data to the correct label) under other poisoning rates on the CIFAR-10 dataset.
+Table VIII and Table IX illustrate the benign accuracy with trigger BadNets and Blended under poisoning rates of and , respectively.
 The results show the similar BA with four strategies, which demonstrates that our proposed strategy has no negative impact on benign accuracy.
@@
 | Avg | 0.886 | 0.886 | 0.885 | 0.885 | 0.877 | 0.877 | 0.875 | 0.876 | 0.883 | 0.883 | 0.885 | 0.885 | |
-A.3 Attack Success Rate Under Additional Target Classes on the CIFAR-10 Dataset
+A-D Attack Success Rate Under Additional Target Classes on the CIFAR-10 Dataset
 This section presents the attack success rate (ASR) under additional target classes on the CIFAR-10 dataset.
-Table 9 illustrates the attack success rate with trigger BadNets and Blended under target category 3.
-Compared with random selection, our proxy-free approach outperforms in all cases, with average boosts of 0.045 and 0.076 on BadNets and Blended triggers.
-The proxy-free method also achieves better attack results in most cases compared to FUS (22/24).
+Table X illustrates the attack success rate with trigger BadNets and Blended under target category 3.
+Compared with random selection, our PFS approach outperforms in all cases, with average boosts of 0.045 and 0.076 on BadNets and Blended triggers.
+The PFS method also achieves better attack results in most cases compared to FUS (22/24).
 The combination of these two achieves superior results in all 24 settings.
@@
 | Avg | 0.796 | 0.848 | 0.842 | 0.885 | 0.816 | 0.916 | 0.859 | 0.933 | |
-A.4 Experiments on Tiny-ImageNet Dataset
-In Table 10, it can be observed that our method outperforms the random method in terms of attack success rates of poisoned samples on Tiny-ImageNet, with average boosts of 0.054, 0.094, and 0.072 on BadNets, Blended, and Optimized triggers, respectively.
-In comparison to FUS, PFS outperforms FUS in 45 out of 54 cases.
-When the two methods are combined, the best results are achieved in 34 out of 54 settings, indicating the superiority of our method.
-These findings provide strong evidence supporting the effectiveness of our proposed method.
+A-E Experiments on CIFAR-100 Dataset
+The results for CIFAR-100 in Table XI are similar to those for CIFAR-10.
+Our PFS approach outperforms random selection in all cases, with average boosts of 0.047, 0.068, and 0.037 for BadNets, Blended, and Optimized triggers, respectively.
+Compared to FUS, PFS achieves better attack results in 49 out of 54 settings.
 | Trigger | Data Transform | Model | |||||||||||
@@
 | Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | Random | PFS | FUS | FUS+PFS | ||
-| BadNets | None | 0.952 | 0.990 | 0.983 | 0.989 | 0.960 | 0.992 | 0.985 | 0.991 | 0.956 | 0.991 | 0.981 | 0.989 |
-| RandomCrop | 0.893 | 0.959 | 0.950 | 0.958 | 0.917 | 0.970 | 0.963 | 0.973 | 0.904 | 0.973 | 0.958 | 0.971 | |
-| RandomHorizontalFlip | 0.952 | 0.989 | 0.984 | 0.989 | 0.963 | 0.991 | 0.981 | 0.989 | 0.947 | 0.983 | 0.974 | 0.985 | |
-| RandomRotation | 0.881 | 0.932 | 0.937 | 0.944 | 0.912 | 0.959 | 0.964 | 0.966 | 0.910 | 0.957 | 0.963 | 0.970 | |
-| ColorJitter | 0.956 | 0.990 | 0.985 | 0.990 | 0.966 | 0.993 | 0.986 | 0.992 | 0.956 | 0.990 | 0.983 | 0.988 | |
-| RandomCrop+ RandomHorizontalFlip | 0.900 | 0.943 | 0.955 | 0.950 | 0.917 | 0.977 | 0.967 | 0.975 | 0.924 | 0.974 | 0.965 | 0.978 | |
-| Avg | 0.922 | 0.967 | 0.966 | 0.970 | 0.939 | 0.980 | 0.974 | 0.981 | 0.933 | 0.978 | 0.971 | 0.980 | |
-| Blended | None | 0.889 | 0.964 | 0.947 | 0.975 | 0.876 | 0.965 | 0.930 | 0.973 | 0.873 | 0.969 | 0.941 | 0.978 |
-| RandomCrop | 0.836 | 0.928 | 0.920 | 0.942 | 0.835 | 0.931 | 0.914 | 0.936 | 0.822 | 0.927 | 0.906 | 0.932 | |
-| RandomHorizontalFlip | 0.905 | 0.966 | 0.964 | 0.980 | 0.888 | 0.964 | 0.950 | 0.977 | 0.886 | 0.965 | 0.954 | 0.981 | |
-| RandomRotation | 0.865 | 0.931 | 0.946 | 0.953 | 0.872 | 0.945 | 0.942 | 0.966 | 0.870 | 0.950 | 0.941 | 0.966 | |
-| ColorJitter | 0.883 | 0.958 | 0.945 | 0.964 | 0.871 | 0.956 | 0.917 | 0.962 | 0.872 | 0.967 | 0.931 | 0.971 | |
-| RandomCrop+ RandomHorizontalFlip | 0.850 | 0.931 | 0.940 | 0.958 | 0.840 | 0.929 | 0.919 | 0.942 | 0.828 | 0.931 | 0.912 | 0.940 | |
-| Avg | 0.871 | 0.946 | 0.944 | 0.962 | 0.864 | 0.948 | 0.929 | 0.959 | 0.859 | 0.952 | 0.931 | 0.961 | |
-| Optimized | None | 0.907 | 0.981 | 0.956 | 0.984 | 0.875 | 0.968 | 0.934 | 0.968 | 0.862 | 0.970 | 0.935 | 0.967 |
-| RandomCrop | 0.897 | 0.976 | 0.959 | 0.980 | 0.886 | 0.968 | 0.950 | 0.972 | 0.869 | 0.958 | 0.945 | 0.951 | |
-| RandomHorizontalFlip | 0.903 | 0.979 | 0.956 | 0.985 | 0.892 | 0.976 | 0.950 | 0.976 | 0.877 | 0.975 | 0.948 | 0.974 | |
-| RandomRotation | 0.899 | 0.977 | 0.957 | 0.972 | 0.897 | 0.978 | 0.956 | 0.972 | 0.888 | 0.973 | 0.948 | 0.956 | |
-| ColorJitter | 0.695 | 0.714 | 0.744 | 0.707 | 0.769 | 0.773 | 0.803 | 0.775 | 0.772 | 0.779 | 0.816 | 0.781 | |
-| RandomCrop+ RandomHorizontalFlip | 0.892 | 0.973 | 0.960 | 0.974 | 0.890 | 0.964 | 0.952 | 0.968 | 0.870 | 0.963 | 0.949 | 0.957 | |
-| Avg | 0.866 | 0.933 | 0.922 | 0.934 | 0.868 | 0.938 | 0.924 | 0.939 | 0.856 | 0.936 | 0.924 | 0.931 | |
-A.5 Experiments on Different Pre-trained Feature Extractor
+| BadNets | None | 0.933 | 0.970 | 0.965 | 0.971 | 0.950 | 0.978 | 0.973 | 0.979 | 0.949 | 0.977 | 0.970 | 0.977 |
+| RandomCrop | 0.882 | 0.947 | 0.930 | 0.945 | 0.894 | 0.951 | 0.940 | 0.952 | 0.888 | 0.951 | 0.940 | 0.951 | |
+| RandomHorizontalFlip | 0.920 | 0.964 | 0.958 | 0.965 | 0.936 | 0.970 | 0.965 | 0.971 | 0.929 | 0.968 | 0.962 | 0.968 | |
+| RandomRotation | 0.868 | 0.923 | 0.912 | 0.924 | 0.866 | 0.939 | 0.925 | 0.938 | 0.875 | 0.935 | 0.926 | 0.938 | |
+| ColorJitter | 0.934 | 0.970 | 0.965 | 0.971 | 0.952 | 0.978 | 0.973 | 0.979 | 0.946 | 0.977 | 0.970 | 0.978 | |
+| RandomCrop+ RandomHorizontalFlip | 0.879 | 0.939 | 0.929 | 0.939 | 0.888 | 0.947 | 0.942 | 0.950 | 0.894 | 0.949 | 0.941 | 0.951 | |
+| Avg | 0.903 | 0.952 | 0.943 | 0.953 | 0.914 | 0.961 | 0.953 | 0.962 | 0.914 | 0.960 | 0.952 | 0.961 | |
+| Blended | None | 0.866 | 0.923 | 0.920 | 0.941 | 0.884 | 0.963 | 0.946 | 0.974 | 0.894 | 0.975 | 0.950 | 0.982 |
+| RandomCrop | 0.863 | 0.874 | 0.891 | 0.880 | 0.826 | 0.916 | 0.906 | 0.938 | 0.829 | 0.931 | 0.912 | 0.942 | |
+| RandomHorizontalFlip | 0.891 | 0.945 | 0.935 | 0.960 | 0.902 | 0.968 | 0.951 | 0.977 | 0.912 | 0.974 | 0.958 | 0.982 | |
+| RandomRotation | 0.818 | 0.862 | 0.901 | 0.865 | 0.841 | 0.916 | 0.916 | 0.936 | 0.848 | 0.926 | 0.921 | 0.944 | |
+| ColorJitter | 0.863 | 0.920 | 0.917 | 0.931 | 0.904 | 0.972 | 0.956 | 0.977 | 0.913 | 0.979 | 0.959 | 0.983 | |
+| RandomCrop+ RandomHorizontalFlip | 0.811 | 0.865 | 0.905 | 0.872 | 0.826 | 0.918 | 0.920 | 0.942 | 0.850 | 0.931 | 0.924 | 0.947 | |
+| Avg | 0.852 | 0.898 | 0.912 | 0.908 | 0.864 | 0.942 | 0.933 | 0.957 | 0.874 | 0.953 | 0.937 | 0.963 | |
+| Optimized | None | 0.979 | 0.999 | 0.995 | 0.999 | 0.884 | 0.972 | 0.931 | 0.982 | 0.895 | 0.983 | 0.962 | 0.987 |
+| RandomCrop | 0.989 | 0.999 | 0.997 | 1.000 | 0.983 | 0.999 | 0.996 | 1.000 | 0.981 | 0.999 | 0.995 | 1.000 | |
+| RandomHorizontalFlip | 0.985 | 0.999 | 0.997 | 1.000 | 0.938 | 0.998 | 0.980 | 0.983 | 0.938 | 0.997 | 0.984 | 0.998 | |
+| RandomRotation | 0.810 | 0.831 | 0.867 | 0.785 | 0.827 | 0.909 | 0.882 | 0.882 | 0.835 | 0.908 | 0.883 | 0.900 | |
+| ColorJitter | 0.984 | 0.997 | 0.993 | 1.000 | 0.967 | 0.989 | 0.979 | 0.985 | 0.956 | 0.993 | 0.990 | 0.994 | |
+| RandomCrop+ RandomHorizontalFlip | 0.991 | 0.999 | 0.997 | 1.000 | 0.987 | 0.999 | 0.998 | 1.000 | 0.982 | 0.998 | 0.994 | 0.999 | |
+| Avg | 0.956 | 0.971 | 0.974 | 0.964 | 0.931 | 0.978 | 0.961 | 0.972 | 0.931 | 0.980 | 0.968 | 0.980 | |
+A-F Experiments on Different Pre-trained Feature Extractor
 To compute the cosine similarity in the proposed PFS, a pre-trained feature extractor is required.